### v2

# The Rees algebra and analytic spread of a divisorial filtration

Abstract.
In this paper we investigate some properties of Rees algebras of divisorial filtrations and their analytic spread. A classical theorem of McAdam shows that the analytic spread of an ideal $I$ in a formally equidimensional local ring is equal to the dimension of the ring if and only if the maximal ideal is an associated prime of $R/\overline{I^{n}}$ for some $n$. We show in Theorem 1.5 that McAdam’s theorem holds for ${\mathbb{Q}}$-divisorial filtrations in an equidimensional local ring which is essentially of finite type over an excellent local ring of dimension less than or equal to 3. This generalizes an earlier result for ${\mathbb{Q}}$-divisorial filtrations in an equicharacteristic zero excellent local domain by the author. This theorem does not hold for more general filtrations.
We consider the question of the asymptotic behavior of the function $n\mapsto\lambda_{R}(R/I_{n})$ for a ${\mathbb{Q}}$-divisorial filtration =$\mathcal{I}=\{I_{n}\}$ of $m_{R}$-primary ideals on a $d$-dimensional normal excellent local ring. It is known from earlier work of the author that the multiplicity
$e(\mathcal{I})=d!\lim_{n\rightarrow\infty}\frac{\lambda_{R}(R/I_{n})}{n^{d}}$
can be irrational. We show in Lemma 4.1 that the limsup of the first difference function
$\limsup_{n\rightarrow\infty}\frac{\lambda_{R}(I_{n}/I_{n+1})}{n^{d-1}}$
is always finite for a ${\mathbb{Q}}$-divisorial filtration. We then give an example in Section 4 showing that this limsup may not exist as a limit.
In the final section, we give an example of a symbolic filtration $\{P^{(n)}\}$ of a prime ideal $P$ in a normal two dimensional excellent local ring which has the property that the set of Rees valuations of all the symbolic powers $P^{(n)}$ of $P$ is infinite.
Key words and phrases:
Rees Algebra, Analytic Spread, Divisorial Valuation, Divisorial Filtration
2020 Mathematics Subject Classification:
13H15, 13A18, 14C17
The first author was partially supported by NSF grant DMS-2054394.
1. Introduction
1.1. Graded Filtrations, Rees Algebras and Analytic Spread
Let $R$ be a Noetherian local ring with maximal ideal $m_{R}$. $\lambda_{R}(M)$ will denote the length of an $R$-module $M$.
If $I$ is an ideal in $R$, then the Rees algebra of $I$ is the graded $R$-algebra $R[It]=\sum_{n\geq 0}I^{n}t^{n}$ and the analytic spread of $I$ is
$\ell(I)=\dim R[It]/m_{R}R[It].$
The classical theory of the analytic spread of an ideal is expounded in [39]. A recent paper on related problems using geometric methods is [34].
It is natural to extend the definition of analytic spread of an ideal to arbitrary graded filtrations of $R$. A graded filtration =$\mathcal{I}=\{I_{n}\}$ of $R$ is a filtration of $R$ such that $I_{m}I_{n}\subset I_{m+n}$ for all $m,n$. Let =$\mathcal{I}=\{I_{n}\}$ be a graded filtration on $R$. The Rees algebra of $\mathcal{I}$ is $R[\mathcal{I}]=\oplus_{n\geq 0}I_{n}$ and the analytic spread of the filtration $\mathcal{I}$ is defined to be
(1) $\ell(\mathcal{I})=\dim R[\mathcal{I}]/m_{R}R[\mathcal{I}].$
Thus if $I$ is an ideal in $R$ and =$\mathcal{I}=\{I^{n}\}$ is the $I$-adic filtration, then $\ell(I)=\ell(\mathcal{I})$.
We always have ([14, Lemma 3.6]) that
(2) $\ell(\mathcal{I})\leq\dim R,$
so the classical inequality $\ell(I)\leq\dim R$ of ideals( [39, Proposition 5.1.6]) extends to arbitrary graded filtrations.
In the case of the $I$-adic filtration =$\mathcal{I}=\{I^{n}\}$ of an ideal in a local ring $R$, the analytic spread $\ell(I)=e$ for some $e$ if and only if $0<\lim_{n\rightarrow\infty}\frac{\lambda_{R}(I^{n}/m_{R}I^{n})}{n^{e-1}}<\infty$. This follows from the theory of Hilbert polynomials of a standard graded ring over a field.
This result fails spectacularly for general filtrations, as $\lambda_{R}(I_{n}/m_{R}I_{n})$ can have arbitrarily large rates of growth on a fixed local ring $R$. This can be seen in the following simple example, which was found with Hailong Dao and Jonathan Montaño.
Example 1.1.
Let $R=k[x,y,z]_{(x,y,z)}$ be the localization of a polynomial ring over a field $k$. Given any function $\sigma:{\mathbb{Z}}_{>0}\rightarrow{\mathbb{Z}}_{>0}$, define the graded filtration =$\mathcal{I}=\{I_{n}\}$ by $I_{0}=R$ and $I_{n}=(z^{n+2})+z^{n+1}(x,y)^{\sigma(n)}$ for $n\geq 1$. Then
$\lambda_{R}(I_{n}/m_{R}I_{n})=\sigma(n)+2.$
The rate of growth of $\lambda_{R}(I_{n}/m_{R}I_{n})$ for “naturally occuring” filtrations is known to be polynomial in many cases. Shankar Dutta in [18] and Hailong Dao and Jonathan Montaño in [17] have shown that the growth of $\lambda_{R}(I^{(n)}/m_{R}I^{(n)})$ for the symbolic filtration $\{I^{(n)}\}$ of an ideal $I$ in a local ring $R$ has polynomial growth if $R$ and the modules $R/I^{(n)}$ have high depth for $n\gg 0$.
When =$\mathcal{I}=\{I_{n}\}$ is an arbitrary graded family of $m_{R}$-primary ideals in a local ring $R$ of dimension $d$, then the rate of growth is always polynomial. We immediately find the bound $\lambda_{R}(I_{n}/m_{R}I_{n})\leq cn^{d}$ for some constant $c$ since there is a power $b$ such that $m_{R}^{b}\subset I_{1}$, so that $m_{R}^{bn}\subset I_{n}$ for all $n$.
1.2. McAdam’s theorem for divisorial filtrations
We have the following remarkable theorem by McAdam.
Theorem 1.2.
([31], [39, Theorem 5.4.6]) Let $R$ be a formally equidimensional local ring and $I$ be an ideal in $R$. Then $m_{R}\in\mbox{Ass}(R/\overline{I^{n}})$ for some $n$ if and only if $\ell(I)=\dim(R)$.
The assumption of being formally equidimensional is not required for the if direction of Theorem 1.2 (this is Burch’s theorem, [4], [39, Proposition 5.4.7 ]).
To see how Theorem 1.2 can be extended, we consider divisorial filtrations. Divisorial filtrations are defined in Section 2.2. Divisorial filtrations arise naturally in ideal theory, as illustrated by the following classical theorem, following from work of Rees.
Theorem 1.3.
Suppose that $I$ is an ideal in an equidimensional excellent local ring $R$. Then the filtration $\{\overline{I^{n}}\}$ of $R$ is a divisorial filtration.
Proof.
This follows from [36] or the proof of [39, Theorem 10.2.2], [39, Proposition 10.4.3] and [21, Scholie IV.7.8.3 (x)]. ∎
The Rees valuations of an ideal $I$ are the divisorial valuations occuring in an irredundant representation of the filtration $\{\overline{I^{n}}\}$ as a divisorial filtration.
Another important example of a divisorial filtration is the filtration of symbolic powers $\{P^{(n)}\}$ of a prime ideal in a local ring $R$ such that $R_{P}$ is a regular local ring. The paper [16] gives a survey of some recent results on symbolic algebras.
The following theorem is proven for local domains in [14]. The proof is the same for local rings, using the notation on valuations of a ring and divisorial filtrations given in Section 2.2.
Theorem 1.4.
(Theorem 1.4 [14] extended to local rings) Suppose that $R$ is a local ring and =$\mathcal{I}=\{I_{n}\}$ is an ${\mathbb{R}}$-divisorial filtration on $R$. Then
  1. a)
    $I_{n}=\overline{I_{n}}$ for all $n>0$ and
  2. b)
    Suppose that $\ell(\mathcal{I})=\dim R$. Then there exists a positive integer $n_{0}$ such that $m_{R}$ is an associated prime of $I_{n}$ for all $n\geq n_{0}$.
The following theorem is a generalization of Theorem 1.2 to divisorial filtrations. If =$\mathcal{I}=\{I_{n}\}$ is a divisorial filtration on an excellent local ring $R$, then $I_{n}=\overline{I_{n}}$ for all $n$ and $R$ is formally equidimensional if $R$ is equidimensional (Scholie IV.7.8.3 (x) [21]).
Theorem 1.5.
Let $R$ be an equidimensional local ring which is essentially of finite type over an excellent local ring of dimension less than or equal to 3. Let =$\mathcal{I}=\{I_{n}\}$ be a ${\mathbb{Q}}$-divisorial filtration on $R$. Then the following are equivalent.
  1. 1)
    The analytic spread of $\mathcal{I}$ is $\ell(\mathcal{I})=\dim R$.
  2. 2)
    There exists $n_{0}\in{\mathbb{Z}}_{>0}$ such that $m_{R}\in\mbox{Ass}(R/I_{n})$ if $n\geq n_{0}$.
  3. 3)
    $m_{R}\in\mbox{Ass}(R/I_{m_{0}})$ for some $m_{0}\in{\mathbb{Z}}_{>0}$.
3) implies 1) of Theorem 1.5 is Theorem 3.5 and Temkin’s theorem on existence of regular alterations [41, Theorem 1.2.5]. Temkin’s theorem shows that if $X$ is an integral scheme which is of finite type over an excellent domain of dimension less than or equal to three, then there exists a projective regular alteration $X^{\prime}\rightarrow X$. Alterations and regular alterations are defined in Definition 3.4 and just after this definition. 1) implies 2) follows from Theorem 1.4.
Theorem 1.5 is proven for excellent equicharacteristic zero local domains in [12, Theorem 1.4], using the existence of resolutions of singularities of excellent equicharacteristic zero domains ([23] or [40]). The proof in [12] shows that Theorem 1.5 is true for excellent local domains of dimension $\leq 3$, using the existence of resolutions of singularities of excellent domains of dimension $\leq 3$ proven in [6].
The conclusions of 3) implies 1) of Theorem 1.5 do not hold for more general filtrations. In fact, there exist ${\mathbb{R}}$-divisorial filtrations $\mathcal{I}$ on an excellent equicharacteristic zero local domain such that the conclusions of Theorem 3.5 are false, as shown in Example 1.5 [12].
The existence of regular alterations over a local domain which is essentially of finite type over a field was first proven by de Jong in [24, Theorem 4.1].
Remark 1.6.
The conclusions of Theorem 1.5 hold whenever $R$ is an equidimensional excellent local ring which has the property that if $P$ is a minimal prime of $R$ and $X\rightarrow\mbox{Spec}(R/P)$ is a birational projective morphism with $X$ integral, then there exists a regular alteration $X^{\prime}\rightarrow X$; in fact, the above proof of Theorem 1.5 (replacing the reference to Temkin’s theorem with this assumption) proves this result.
1.3. Hilbert functions of divisorial $m_{R}$-filtrations
If $I$ is an ideal in a Noetherian local ring $R$, the function $\lambda_{R}(I^{n}/m_{R}I^{n})$ is a polynomial for large $n$, whose degree is the analytic spread $\ell(I)$ of $I$. The computation of asymptotic properties of the function $\lambda_{R}(I_{n}/m_{R}I_{n})$ is difficult for a divisorial filtration =$\mathcal{I}=\{I_{n}\}$. In the case that $R$ is normal and excellent of dimension two, we show in [11, Theorem 1.6] that for a divisorial filtration $\mathcal{I}$ of $R$, there exists $\alpha\in{\mathbb{Q}}_{\geq 0}$ such that
$\lambda_{R}(I_{n}/m_{R}I_{n})=n\alpha+\sigma(n)$
for $n>0$, where $\sigma(n)$ is a bounded function of $n$. We have that $\alpha>0$ if and only if $\ell(\mathcal{I})=2=\dim R$. The proof uses Zariski decomposition, which does not exist in general if $\dim R\geq 3$.
Now suppose that $\mathcal{I}$ is a divisorial filtration of $m_{R}$ ideals. Then, related to the function $\lambda_{R}(I_{n}/m_{R}I_{n})$, we have the functions
(3) $n\mapsto\lambda_{R}(R/I_{n}).$
and its first difference $n\mapsto\lambda_{R}(I_{n}/I_{n+1})$. A good property of the function (3) is that the multiplicity
$e(\mathcal{I})=\lim_{n\rightarrow\infty}d!\frac{\lambda_{R}(R/I_{n})}{n^{d}}$
always exists. In fact, this limit always exists for arbitrary graded families of $m_{R}$-primary ideals in analytically unramified local rings ([7, Theorem 1.1] and [8, Theorem 4.2]). This theorem builds on earlier work by [19], [32], [33], [25] and [28].
The multiplicity $e(\mathcal{I})$ is related to the analytic spread $\ell(\mathcal{I})$ as follows.
Lemma 1.7.
Suppose that $R$ is a $d$-dimensional excellent normal local ring and =$\mathcal{I}=\{I_{n}\}$ is a divisorial filtration of $m_{R}$-primary ideals of $R$. Then $e(\mathcal{I})>0$ and $\ell(\mathcal{I})=d$.
We give an outline of the proof. $\ell(\mathcal{I})=d$ by Theorem 1.5. There exist divisorial valuations $w_{j}$ on $R$ (the Rees valuations of $m_{R}$) such that $\mbox{center}_{R}(w_{j})=m_{R}$ for all $j$, and positive integers $b_{j}$ such that
$\overline{m_{R}^{n}}=\cap_{j}I(w_{j})_{mb_{j}}\mbox{ for all $n\geq 0$}.$
Writing
$I_{n}=\cap_{j}I(v_{i})_{na_{i}}$
for suitable divisorial valuations $v_{i}$ with $\mbox{center}_{R}(v_{i})=m_{R}$ and $a_{i}\in{\mathbb{Q}}$, we see by Izumi’s Theorem ([37]) that there exists a positive integer $\beta$ such that $I_{\beta n}\subset\overline{m_{R}^{n}}$ for all $n>0$. Now the ring $\sum_{n\geq 0}\overline{m_{R}^{n}}t^{n}$ is a finite $R[m_{R}t]$-module by [39, Proposition 5.2.1] and [39, Corollary 9.2.1], so
$0<\lim_{n\rightarrow\infty}\frac{\lambda_{R}(R/\overline{m_{R}^{n}})}{n^{d}}\leq\lim_{n\rightarrow\infty}\frac{\lambda_{R}(R/I_{n})}{n^{d}}=e(\mathcal{I}).$
In this subsection, we consider the asymptotic properties of the difference function $n\mapsto\ell_{R}(I_{n}/I_{n+1})$. In the case of an $I$-adic filtration, this function is a polynomial of degree $d-1$ for $n\gg 0$. We always have that
(4) $\limsup_{n\rightarrow\infty}\frac{\lambda_{R}(I_{n}/I_{n+1})}{n^{d-1}}$
is finite for ${\mathbb{Q}}$-divisorial filtrations of $m_{R}$-primary ideals, as is shown in Lemma 4.1. However, there are examples of graded families of ideals such that this limsup is infinite ([8, Theorem 4.3]). Even when this limsup is finite, it may be that the limsup is not a limit for graded filtrations of $m_{R}$-primary ideals ([8, Theorem 4.6]).
The function (3) is very nice in dimension 2. In Section 8 [15], it is shown that if $R$ is a two dimensional normal excellent local ring and =$\mathcal{I}=\{I_{n}\}$ is a ${\mathbb{Q}}$-divisorial filtration of $m_{R}$-primary ideals on $R$, then there exists a quadratic polynomial $P(n)$ with rational coefficients, a function $f(n)$ and a constant $c$ such that
$\lambda_{R}(R/I_{n})=P(n)+f(n)$
with $|f(n)|<c$ for all $n$. In [15, Theorem 9], it is shown that if $R$ has equicharacteristic zero then $f(n)$ is eventually periodic. Example 5 [15] shows that $f(n)$ may not be eventually periodic if $R$ has characteristic $p>0$.
It follows that in dimension two, the multiplicity $e(\mathcal{I})$ is always a rational number. Further, the limsup (4) is a limit, and this limit is a rational number.
In Example 6 [15] and in [10], examples are given of a three dimensional normal local domain $R$ with an isolated singularity and a divisorial valuation $v$ on $R$ such that the center of $v$ on $R$ is the maximal ideal $m_{R}$ and the multiplicity $e(\mathcal{I})$ of the filtration =$\mathcal{I}=\{I_{n}\}$ where $I_{n}=\{x\in R\mid v(x)\geq n\}$ is irrational.
By (4), the limsup of difference functions always exists. We consider in Section 4 the question of if this limsup is always a limit; that is, does the limit
(5) $\lim_{n\rightarrow\infty}\frac{\lambda_{R}(I_{n}/I_{n+1})}{n^{d-1}}$
of a ${\mathbb{Q}}$-divisorial filtration of $m_{R}$-primary ideals always exist?
As we have seen above, this limit does always exist and is even a rational number when the dimension of $R$ is $2$.
We give in Section 4 an example showing that the limit (5) does not exist for a ${\mathbb{Q}}$-divisorial filtration on a three dimensional normal local ring $R$ which is essentially of finite type over an algebraically closed field of characteristic zero.
To accomplish this, we start with the construction of an example with irrational multiplicity in [10]. In this example, a resolution of singularities $U\rightarrow\mbox{Spec}(R)$ of $\mbox{Spec}(R)$ is constructed which has two exceptional divisors, $\overline{S}$ and $F$ both of which contract to $m_{R}$. Let $v$ be the canonical valuation associated to the valuation ring $\mathcal{O}_{F}$ and =$\mathcal{I}=\{I_{n}\}$ where $I_{n}=I(v)_{n}=\{x\in R\mid v(x)\geq n\}$, which is a ${\mathbb{Z}}$-divisorial filtration on $R$.
It is shown in the formula before [10, Lemma 2.2] and [10, Theorem 4.1] that
$I_{n}=I(nF):=\Gamma(U,\mathcal{O}_{U}(-nF))=\Gamma(U,\mathcal{O}_{U}(-D_{n}))$
where $D_{n}=\lceil n\alpha\rceil\overline{S}+nF$, $\alpha=\frac{3}{9-\sqrt{3}}$. It is further shown that the divisors $-D_{n}$ are nef divisors on $U$. Thus by formula (9) and [10, Proposition 2.4],
$\lim_{n\rightarrow\infty}\frac{\lambda_{R}(R/I(nF))}{n^{3}}=-\frac{((-\alpha\overline{S}-F)^{3})}{3!}=\alpha^{3}468-486\alpha^{2}+162\alpha+54,$
so that
(6) $e(\mathcal{I})=3!(\alpha^{3}468-486\alpha^{2}+162\alpha+54)=\frac{72252}{169}-\frac{162}{169}\sqrt{3}$
is an irrational number. Equation (6) corrects equation (12) of [10], where $\alpha^{3}468-486\alpha^{2}+162\alpha+54$ is incorrectly calculated to be $\frac{12042}{169}-\frac{27}{169}\sqrt{3}$.
To compute the second difference $\lambda_{R}(I(nF)/I(n+1)F)$, we must make a finer calculation, computing the function $n\mapsto n\lambda_{R}(R/I_{n})$ explicitely up to a constant times $n$.
In (39) it is shown that
(7) $\lambda_{R}(R/I(nF))=\frac{1}{6}(D_{n}^{3})+\frac{1}{4}(D_{n}^{2}\cdot K_{Y})+\delta(n)$
for $n\in{\mathbb{N}}$ where there is a constant $c_{1}$ such that $|\delta(n)|<c_{1}n$ for all $n\in{\mathbb{N}}$. By (41) and (45), this length has the numeric formulation
$\begin{array}[]{lll}\lambda_{R}(R/I(nF))&=&\frac{486}{6}\lceil\alpha n\rceil^{3}-\frac{486}{6}n\lceil\alpha n\rceil^{2}+\frac{162}{6}n^{2}\lceil\alpha n\rceil+\frac{54}{6}n^{3}\\ &&-\frac{792}{4}\lceil n\alpha\rceil^{2}+\frac{564}{4}n\lceil n\alpha\rceil-\frac{174}{4}n^{2}+\delta(n).\end{array}$
So taking the limit in this equation, we obtain
(8) $e(\mathcal{I}(F))=3!\left(\lim_{n\rightarrow\infty}\frac{\lambda_{R}(R/I(nF))}{n^{3}}\right)=3!(\alpha^{3}468-486\alpha^{2}+162\alpha+54)=\frac{72252}{169}-\frac{162}{169}\sqrt{3}$
obtaining (6) by a different method.
From (7), we calculate in (46) that
(9) $\lambda_{R}(I(nF)/I((n+1)F))=\lambda_{R}(R/I((n+1)F))-\lambda_{R}(R/I(nF))=\frac{1}{6}\left((D_{n+1}^{3})-(D_{n}^{3})\right)+\varepsilon(n)$
where there is a constant $c^{\prime\prime}$ such that $|\varepsilon(n)|<c^{\prime\prime}n$ for $n\in{\mathbb{N}}$. At the end of Subsection 4.5, it is shown that ${\mathbb{Z}}_{>0}$ is the union ${\mathbb{Z}}_{>0}=\Sigma_{1}\coprod\Sigma_{2}$ where both $\Sigma_{1}$ and $\Sigma_{2}$ are infinite, the limit as $n\rightarrow\infty$ restricted to $n\in\Sigma_{1}$ of $\frac{\lambda_{R}(I(nF)/I((n+1)F)}{n^{2}}$ is
$\frac{1}{6}\left(-486\alpha^{2}+324\alpha+162\right)=\frac{144504}{6(26)^{2}}-\frac{324}{6(26)^{2}}\sqrt{3}$
and the limit as $n\rightarrow\infty$ restricted to $n\in\Sigma_{2}$ of $\frac{\lambda_{R}(I(nF)/I((n+1)F)}{n^{2}}$ is
$\frac{1}{6}\left(918\alpha^{2}-810\alpha+324\right)=\frac{106596}{6(26)^{2}}-\frac{4536}{6(26)^{2}}\sqrt{3}.$
These two limits, over $n$ in $\Sigma_{1}$ and $\Sigma_{2}$ respectively are not equal, so the limit
$\lim_{n\rightarrow\infty}\frac{\lambda_{R}(I_{n}/I_{n+1})}{n^{2}}$
does not exist.
1.4. Symbolic powers with infinitely many Rees valuations
Let $R$ be a $d$-dimensional excellent normal local ring. Let =$\mathcal{I}=\{I_{n}\}$ be a graded family of ideals of $R$ ($\mathcal{I}$ is a filtration of $R$ and $I_{m}I_{n}\subset I_{m+n}$ for all $m,n$). We have that $\sqrt{I_{n}}=\sqrt{I_{1}}$ for all $n$. We define the Rees valuations of $\mathcal{I}$ to be the divisorial valuations $v$ of $R$ such that the center of $v$ on $\mbox{Proj}(R[\mathcal{I}])$ has dimension $d-1$ and the center $\mbox{center}_{R}(v)$ of $v$ on $R$ contains $I_{1}$ (and so contains $I_{n}$ for all $n$). In the case that =$\mathcal{I}=\{I^{n}\}$ is the $I$-adic filtration for some ideal $I$ of $R$, this agrees with the classical definition of the Rees valuations of $I$ ([39, Chapter 10]).
Suppose that $\mathcal{I}$ is a divisorial filtration with $I_{n}=I(v_{1})_{n\lambda_{1}}\cap\cdots\cap I(v_{r})_{n\lambda_{r}}$ for all $n$. Then the Rees valuations of $\mathcal{I}$ are a subset of $\{v_{1},\ldots,v_{r}\}$. The Rees valuations can be a proper subset of $\{v_{1},\ldots,v_{r}\}$. Theorem 1.5 shows that if $R$ is essentially of finite type over a field (in addition to our assumption that $R$ is normal and excellent) then the analytic spread $\ell(\mathcal{I})=d$ if and only if there exists a Rees valuation $v$ of $\mathcal{I}$ such that $\mbox{center}_{R}(v)=m_{R}$.
We have the natural birational maps
$\mbox{Proj}(R[I_{n}t^{n}])\dashrightarrow\mbox{Proj}(R[\mathcal{I}])$
for all $n>0$. The nature of these birational maps (and hence the structure of $\mbox{Proj}(R[\mathcal{I}])$) is determined by the respective Rees valuations of $\mbox{Proj}(R[I_{n}t^{n}])$ and $\mbox{Proj}(R[\mathcal{I}])$.
This raises the question of what asymptotic properties the Rees valuations of $I_{n}$ have (as $n$ goes to infinity) and how this compares with the Rees valuations of $\mathcal{I}$.
If $I$ is an ideal in an excellent local ring, then we have seen in Theorem 1.3 that the filtration of powers of integral closures $\{\overline{I^{n}}\}$ of the powers of $I$ is a ${\mathbb{Z}}$-divisorial filtration. In particular, the set of all Rees valuations of all of the ideals $I^{m}$ over all $m$ is a finite set. It is just the set of Rees valuations of $I$ itself. In Section 5 we consider the question of if this set is finite for the ideals in a ${\mathbb{Z}}$-divisorial filtration; that is, is the set of all Rees valuations of the ideals $I_{n}$ finite when =$\mathcal{I}=\{I_{n}\}$ is a ${\mathbb{Z}}$-divisorial filtration? We show that this is not the case. In Section 5 we construct an example of a height one prime ideal $P$ in a two dimensional normal local ring $R$ such that for all $n\geq 1$, the symbolic power $P^{(n)}$ has a Rees valuation other then the $PR_{P}$-adic valuation $v_{P}$, but for $1\leq m<n$, $v_{P}$ is the only common Rees valuation of $P^{(m)}$ and $P^{(n)}$. In particular, the set of Rees valuations of $P^{(n)}$ over all $n\geq 1$ is infinite.
The only Rees valuation of the divisorial filtration $\{P^{(n)}\}$ is $v_{P}$ and the analytic spread of this filtration is ($\ell(\{P^{(n)}\})=0$. In contrast, the analytic spread of each individual ideal $P^{(n)}$ is maximal, with $\ell(P^{(n)})=2$. None of the Rees valuations of the $P^{(n)}$ which have center $m_{R}$ on $R$ have a center on $\mbox{Proj}(R[\mathcal{I}])=\mbox{Spec}(R)\setminus\{m_{R}\}$. Necessarily, $R[\mathcal{I}]$ is not Noetherian.
1.5. Acknowledgements
The author would like to thank Hailong Dao and Jonathan Montaño for discussions and questions about Hilbert polynomials of filtrations and to thank Linquin Ma, Thomas Polstra and Karl Schwede for a discussion about the cardinality (finite or infinite) of the set of Rees valuations of the symbolic powers of a prime ideal. He thanks Suprajo Das for suggesting the use of alterations to generalize the assumption that rings have equicharacteristic zero in [12, Theorem 1.3]. He thanks the reviewer for helpful comments to improve the exposition.
2. Valuations, Filtrations and Rees Algebras
A local ring will be assumed to be Noetherian. We will denote the maximal ideal of a local ring $R$ by $m_{R}$. The length of an $R$-module $M$ will be denoted by $\lambda_{R}(M)$ or $\lambda(M)$.
2.1. Extensions of valuations and Rees algebras
Let $K$ be a field and $v$ be a rank 1 valuation of $K$ (the value group $vK$ of $v$ is order isomorphic to a subgroup of ${\mathbb{R}}$). For $\lambda\geq 0$ in the value group $vK$, let
$I(v)_{\lambda}=\{x\in K\mid v(c)\geq\lambda\},$
an ideal in the valuation ring $\mathcal{O}_{v}$ of $v$. Let $R$ be a domain with quotient field $K$ such that $v$ is nonnegative on $R$. The center of $v$ on $R$ is the prime ideal
$\mbox{center}_{R}(v)=m_{v}\cap R,$
where $m_{v}$ is the maximal ideal of $\mathcal{O}_{v}$. Let $P=\mbox{center}_{R}(v)$. We have that $\sqrt{I(v)_{\lambda}\cap R}=P$ for all $\lambda\in vK_{>0}$ and in fact $I(v)_{\lambda}\cap R$ is $P$-primary, since the inclusion $R\rightarrow\mathcal{O}_{v}$ factors through $R_{P}$ and $m_{v}\cap R_{P}=PR_{P}$.
The following lemma is immediate.
Lemma 2.1.
Let $K\rightarrow L$ be a finite extension of fields, $v_{1},\ldots,v_{r}$ be rank 1 valuations of $K$, $\lambda_{i}\in v_{i}K$ and $\{w_{ij}\}$ be the extensions of $v_{i}$ to $L$ for $1\leq i\leq r$. Then
$(\cap_{i}\cap_{j}I(w_{ij})_{\lambda_{i}})\cap K=\cap_{i}I(v_{i})_{\lambda_{i}}.$
Lemma 2.2.
Let $K\rightarrow L$ be a finite field extension, $R$ be a normal Noetherian domain with quotient field $K$ and $S$ be the integral closure of $R$ in $L$. Let $v$ be a valuation of $K$ which is nonnegative on $R$, and $w$ be an extension of $v$ to $L$. Then $w$ is nonnegative on $S$ and
$\dim S/\mbox{center}_{S}(w)=\dim R/\mbox{center}_{R}(v).$
In particular, $\mbox{center}_{R}(v)$ is a maximal ideal of $R$ if and only of $\mbox{center}_{S}(w)$ is a maximal ideal of $S$.
Proof.
The maximal ideal $m_{w}$ of $\mathcal{O}_{w}$ contracts to the maximal ideal $m_{v}$ of $\mathcal{O}_{v}$ so $\mbox{center}_{S}(w)$ contracts to $\mbox{center}_{R}(v)$. Thus the lemma holds since $S$ is integral over $R$. ∎
For ideals $U$ and $V$ of a ring $R$, define
$U:V^{\infty}=\cup_{n>0}U:V^{n}=\{f\in R\mid fV^{n}\subset U\mbox{ for some $n>0$}\}.$
If $R$ is Noetherian, $U:V^{\infty}$ is obtained from $U$ by removing from a primary decomposition of $U$ the $P$-primary components that have the property that $J(R)\subset P$.
Let $R$ be a normal Noetherian semi local domain with quotient field $K$, $v_{1},\ldots,v_{r}$ be rank 1 valuations of $K$ which are nonnegative on $R$ and $\lambda_{i}\in v_{i}K_{\geq 0}$ for $1\leq i\leq r$. Let $J(R)$ be the intersection of the maximal ideals of $R$.
We have that
$\left(\cap_{i=1}^{r}I(v_{i})_{\lambda_{i}}\right)\cap R=\cap_{i=1}^{r}(I(v_{i})_{\lambda_{i}}\cap R).$
Let $P_{j}$ for $1\leq j\leq s$ be the distinct centers of the $v_{i}$ on $R$, and define $\sigma(i)$ by $\mbox{center}_{R}(v_{i})=P_{\sigma(i)}$ for $1\leq i\leq r$. As observed before Lemma 2.1, $I(v_{i})_{\lambda_{i}}\cap R$ is $P_{\sigma(i)}$-primary for all $i$. Let $Q_{j}=\cap_{\sigma(i)=j}(I(v_{i})_{\lambda_{i}}\cap R)$, which is a $P_{j}$-primary ideal since it is an intersection of $P_{j}$-primary ideals. Thus $\cap_{j=1}^{s}Q_{j}$ is a primary decomposition of $\cap_{i=1}^{r}(I(v_{i})_{\lambda_{i}}\cap R)$. As remarked above, $(\cap_{j=1}^{s}Q_{j}):J(R)^{\infty}$ removes from $\cap_{j=1}^{s}$ the primary components from the maximal ideals of $R$.
Thus
(10) $\left[\cap_{i\in\Lambda}I(v_{i})_{\lambda_{i}}\right]\cap R=\left(\left[\cap_{i=1}^{r}I(v_{i})_{\lambda_{i}}\right]\right)\cap R:J(R)^{\infty},$
where
$\Lambda=\{i\mid\mbox{center}_{R}(v_{i})\mbox{ is not a maximal ideal of $R$}\}.$
The following lemma follows from Lemma 2.1, Lemma 2.2 and equation (10).
Lemma 2.3.
Let $K\rightarrow L$ be a finite field extension, $R$ be a normal Noetherian domain with quotient field $K$ and $S$ be the integral closure of $R$ in $L$. Suppose that $S$ is a finite $R$-module. Let $v_{1},\ldots,v_{r}$ be rank 1 valuations of $K$, $\lambda_{i}\in v_{i}K$ be nonnegative and $\{w_{ij}\}$ be the extensions of $v_{i}$ to $L$ for $1\leq i\leq r$. Suppose that $v_{1},\ldots,v_{r}$ are nonnegative on $R$. Then
$\left(\left[\cap I(w_{ij})_{\lambda_{i}}\right]\cap S:J(S)^{\infty}\right)\cap R=\left[\cap I(v_{i})_{\lambda_{i}}\right]\cap R:J(R)^{\infty}.$
In particular,
$\left[\cap I(v_{i})_{\lambda_{i}}\right]\cap R:J(R)^{\infty}\neq\left[\cap I(v_{i})_{\lambda_{i}}\right]\cap R$
implies
$\left[\cap I(w_{ij})_{\lambda_{i}}\right]\cap S:J(S)^{\infty}\neq\left[\cap I(w_{ij})_{\lambda_{i}}\right]\cap S.$
Let $K$ be a field and $v_{1},\ldots,v_{r}$ be rank 1 valuations of $K$ and $\lambda_{i}\geq 0$ be elements of $v_{i}K$ for $1\leq i\leq r$. Let $R$ be a normal semilocal domain with quotient field $K$ such that the $v_{i}$ are nonnegative on $R$ for all $i$. Define a graded filtration $\mathcal{I}(R)=\{I(R)_{n}\}$ of $R$ by $I(R)_{n}=(\cap_{i}I(v_{i})_{n\lambda_{i}})\cap R$. Suppose that $K\rightarrow L$ is a finite field extension. Let $\{w_{ij}\}$ be the extensions of $v_{i}$ to $L$ for $1\leq i\leq r$. Let $S$ be the integral closure of $R$ in $L$. Define a graded filtration
(11) $\mathcal{I}(S)=\{I(S)_{n}\}$ on $S$ by $I(S)_{n}=(\cap_{i}\cap_{j}I(w_{ij})_{n\lambda_{i}})\cap S$.
Proposition 2.4.
Let $L$ be a finite field extension of $K$ and $S$ be the integral closure of $R$ in $L$. Then $S[\mathcal{I}(S)]$ is integral over $R[\mathcal{I}(R)]$.
Proof.
Let $N$ be a finite extension of $L$ such that $N$ is a normal extension of $K$. Let $G=\mbox{Aut}_{K}(N)$. We have a tower of fields $K\rightarrow N^{G}\rightarrow N$ where $N^{G}$ is the fixed field of $N$ under the action of $G$. The extension $K\rightarrow N^{G}$ is purely inseparable and $N^{G}\rightarrow N$ is Galois with Galois group $G$ (c.f. [27, Proposition V.6.11]).
Let $A$ be the integral closure of $R$ in $N^{G}$ and $B$ be the integral closure of $R$ in $N$. To prove the proposition, it suffices to show that $B[\mathcal{I}(B)]$ is integral over $R[\mathcal{I}(R)]$, and to prove this it suffices to prove that $B[\mathcal{I}(B)]$ is integral over $A[\mathcal{I}(A)]$ and $A[\mathcal{I}(A)]$ is integral over $R[\mathcal{I}(R)]$. Thus we are reduced to proving the Proposition in two cases: when $L$ is Galois over $K$ and when $L$ is purely inseparable over $K$.
First suppose that $L$ is Galois over $K$, with Galois group $G$. If $x\in S^{G}$, then $x\in K$, and since $x$ is integral over $R$, $x\in R$. Thus $R=S^{G}$. Let
$G=\{\sigma_{1}={\rm id},\sigma_{2},\ldots,\sigma_{a}\}.$
Let $x\in S$. Let $f(X)=\prod_{i=1}^{a}(X-\sigma_{i}(x))\in K[X]$. We have that $f(x)=0$. Expand
(12) $f(X)=X^{a}-s_{1}X^{a-1}+s_{2}X^{a-2}+\cdots+(-1)^{a}s_{a}$
where $s_{i}$ are the elementary symmetric polynomials in $\sigma_{1}(x),\ldots,\sigma_{a}(x)$. The $s_{i}$ are integral over $R$, and are contained in $K$, so that the $s_{i}$ are in $R$ since $R$ is normal.
Suppose that $x\in I(S)_{n}\subset S$. Then
(13) $w_{i,j}(x)\geq n\lambda_{i}$
for all $i,j$. There exists a permutation $\tau_{i}$ of $\{1,\ldots,r\}$ such that
(14) $w_{i,1}\sigma_{j}=w_{i,\tau_{i}(j)}$
for all $j$ ([42, Corollary VI.7.3]). For $b\in\{1,\ldots,a\}$, we compute
$v_{i}(s_{b})=w_{i,1}(s_{b})\geq\min\{w_{i,1}(\sigma_{\alpha_{1}}(x))+\cdots+w_{i,1}(\sigma_{\alpha_{b}}(x))\}$
where the minimum is over $\alpha_{1}<\alpha_{2}<\cdots<\alpha_{b}$. Thus
(15) $v_{i}(s_{b})\geq bn\lambda_{i}$
by (13) and (14), so that $s_{b}\in I(R)_{bn}$.
Write $R[\mathcal{I}(R)]\cong\sum_{i=0}^{\infty}I(R)_{n}t^{n}$, a graded subring of the graded polynomial ring $R[t]$. Let $g(X)=\prod_{i=1}^{a}(X-\sigma_{i}(x)t^{n})$. Expand
$g(X)=X^{a}-s_{1}t^{n}X^{a-1}+s_{2}t^{2n}X^{ra2}+\cdots+(-1)^{a}s_{a}t^{an}.$
We have that $(-1)^{b}s_{b}t^{bn}\in I(R)_{bn}t^{bn}$ for all $b$ by (15). Thus $g(X)\in R[\mathcal{I}(R)][X]$. Since
$g(xt^{n})=\prod_{i=1}^{a}(xt^{n}-\sigma_{i}(x)t^{n})=t^{an}\prod_{i=1}^{a}(x-\sigma_{i}(x))=0,$
this is an equation of integral dependence of $xt^{n}$ over $R[\mathcal{I}(R)]$. Thus $S[\mathcal{I}(S)]$ is integral over $R[\mathcal{I}(R)]$ in the case that $L$ is Galois over $K$.
Now suppose that $L$ is purely inseparable over $K$. Let $[L:K]=p^{m}$ where $p$ is the characteristic of $K$. If $x\in L$, we have that $x^{p^{m}}\in K$. There is a unique extension $w_{i}$ of $v_{i}$ to $L$. This valuation satisfies $w_{i}(x)=\frac{1}{p^{m}}v_{i}(x^{p^{m}})$ for $x\in L$. Thus $x\in I(S)_{n}$ implies $x^{p^{m}}\in I(R)_{np^{m}}$ and so $(xt^{n})^{p^{m}}\in I(R)_{np^{m}}t^{np^{m}}$. Let $g(X)=X^{p^{m}}-(xt^{n})^{p^{m}}\in R[\mathcal{I}(R)][X]$. We have that $g(xt^{n})=0$, giving an equation of integral dependence of $xt^{n}$ over $R[\mathcal{I}(R)]$. Thus $S[\mathcal{I}(S)]$ is integral over $R[\mathcal{I}(R)]$ in the case that $L$ is purely inseparable over $K$.
2.2. Divisorial Filtrations
A divisorial valuation of a Noetherian domain $R$ ([39, Definition 9.3.1]) is a valuation $\nu$ of $K$ such that if $\mathcal{O}_{\nu}$ is the valuation ring of $\nu$ with maximal ideal $\mathfrak{m}_{\nu}$, then $R\subset\mathcal{O}_{\nu}$ and if $\mathfrak{p}=\mathfrak{m}_{\nu}\cap R$ then $\mbox{trdeg}_{\varkappa(\mathfrak{p})}\varkappa(\nu)={\rm ht}(\mathfrak{p})-1$, where $\varkappa(\mathfrak{p})$ is the residue field of $R_{\mathfrak{p}}$ and $\varkappa(\nu)$ is the residue field of $\mathcal{O}_{\nu}$. If $R$ is a local ring and $\nu$ is a divisorial valuation of $R$ such that $m_{R}=m_{\nu}\cap R$, then $\nu$ is called an $m_{R}$-valuation.
By [39, Theorem 9.3.2], the valuation ring of every divisorial valuation $\nu$ is Noetherian, hence is a discrete valuation. Suppose that $R$ is an excellent local domain. Then a valuation $\nu$ of the quotient field $K$ of $R$ which is nonnegative on $R$ is a divisorial valuation of $R$ if and only if the valuation ring $\mathcal{O}_{\nu}$ is essentially of finite type over $R$ ([13, Lemma 6.1] or [39, Theorem 9.3.2]).
Let $R$ be a Noetherian domain and $v$ be a discrete valuation of the quotient field of $R$ which is nonnegative on $R$. For $\lambda\in{\mathbb{R}}_{\geq 0}$, define the valuation ideal of $R$
$I_{R}(v)_{\lambda}=\{f\in R\mid v(f)\geq\lambda\}=I(v)_{\lambda}\cap R.$
A divisorial filtration of a Noetherian domain $R$ is a filtration =$\mathcal{I}=\{I_{n}\}$ where there exist divisorial valuations $v_{1},\ldots,v_{r}$ of $R$ and $\lambda_{1},\ldots,\lambda_{r}\in{\mathbb{R}}_{\geq 0}$ such that
$I_{n}=I_{R}(v_{1})_{n\lambda_{1}}\cap\cdots\cap I_{R}(v_{r})_{n\lambda_{r}}=I(v_{1})_{n\lambda_{1}}\cap\cdots\cap I(v_{r})_{n\lambda_{r}}\cap R.$
We now weaken our assumptions on $R$ and suppose that $R$ is an arbitrary Noetherian local ring.
Definition 2.5.
Let $R$ be a Noetherian local ring and $P_{1},\ldots,P_{t}$ be the minimal prime ideals of $R$. Let $\varphi_{i}:R\rightarrow R/P_{i}$ be the natural projection for $1\leq i\leq t$. Let $\{v_{ij}\}$ be divisorial valuations on $R/P_{i}$ and $\lambda_{ij}\in{\mathbb{R}}_{\geq 0}$. Let =$\mathcal{I}=\{I_{n}\}$ where
(16) $I_{n}=\cap_{i=1}^{t}\cap_{j}\varphi_{i}^{-1}(I_{R/P_{i}}(v_{ij})_{n\lambda_{ij}}).$
$\mathcal{I}$ is a graded filtration of $R$, which is called an ${\mathbb{R}}$-divisorial filtration. If the $\lambda_{ij}$ are in ${\mathbb{Q}}_{\geq 0}$ then $\mathcal{I}$ is called a ${\mathbb{Q}}$-divisorial filtration and if the $\lambda_{ij}$ are in ${\mathbb{Z}}_{\geq 0}$ then $\mathcal{I}$ is called a divisorial filtration.
If $P$ is a prime ideal of a local ring $R$, $\varphi:R\rightarrow R/P$ and $v$ is a valuation of the quotient field of $R/P$, then $w=v\circ\varphi$ is a valuation of $R$, as defined in VI.3.1 [2]. That is, by extending the value group $vR/P$ of $V$ by introducing the element $\infty$, which is larger than any element of $vR/P$, and letting $\Gamma_{\infty}=vR/P\cup\{\infty\}$, the mapping $w:R\rightarrow\Gamma_{\infty}$ satisfies the properties
(17) $\begin{array}[]{l}w(xy)=w(x)+w(y)\mbox{ for }x,y\in R\\ w(x+y)\geq\min\{w(x),w(y)\}\mbox{ for }x,y\in R\\ w(1)=0\mbox{ and }w(P)=\infty\end{array}$
With the notation of (16), the mappings $v_{ij}\circ\varphi_{i}$ are valuations of $R$, satisfying the conclusions of (17).
Associated to a discrete valuation ring $V$ is a unique canonical valuation $v_{V}$ defined by $v_{V}(x)=n$ if $x\in m_{V}^{n}\setminus m_{V}^{n+1}$ for $x\in V$ and where $m_{V}$ is the maximal ideal of $V$. If $w$ is any valuation whose valuation ring is $V$, so that $w$ is an equivalent valuation to $v$, we have that $w=cv$ where $c=w(u)$ for $u$ a generator of $m_{V}$ (a uniformizing parameter of $V$).
Now suppose that $R$ is a normal excellent local domain. Let
(18) $\pi:X\rightarrow\mbox{Spec}(R)$
be a birational projective morphism such that $X$ is normal. Let
(19) $D=\sum\lambda_{i}E_{i}$
be an effective ${\mathbb{R}}$-Weil divisor on $X$ (so that all $\lambda_{i}\in{\mathbb{R}}_{\geq 0}$), where the $E_{i}$ are prime divisors on $X$. Define
$\lceil nD\rceil=\sum\lceil n\lambda_{i}\rceil E_{i},$
where $\lceil x\rceil$ is the round up of a real number $x$. Then for all $n$, $I(nD)=\Gamma(X,\mathcal{O}_{X}(-\lceil nD\rceil))$ is an ideal of $R$ and =$\mathcal{I}=\{I(nD)\}$ is a divisorial filtration of $R$.
Let $v_{1},\ldots,v_{r}$ be the canonical valuations of the valuation rings $\mathcal{O}_{X,E_{i}}$. Then
(20) $\Gamma(X,\mathcal{O}_{X}(-\lceil nD\rceil))=I_{R}(v_{1})_{n\lambda_{1}}\cap\cdots\cap I_{R}(v_{r})_{n\lambda_{r}}.$
Let =$\mathcal{I}=\{I_{n}\}$ be a divisorial filtration of $R$. By [13, Remark 6.6 to Lemma 5.5], there exists $\pi:X\rightarrow\mbox{Spec}(R)$ and $D$ as in (18) and (19) such that $I_{n}=I(nD)$ for all $n\geq 0$.
A choice of $X$ and $D$ such that $I_{n}=I(nD)$ for all $n\geq 0$ is called a representation of the divisorial filtration $\mathcal{I}$. The choice of $X$ is not unique, and the choice of $D$ on a particular $X$ may not be unique. It is possible for a particular $\mathcal{I}$ to have representations (20) which are ${\mathbb{Q}}$-divisorial and also have representations (20) which are not ${\mathbb{Q}}$-divisorial.
Let $K$ be the quotient field of $R$, $K\rightarrow L$ be a finite extension of fields and $\{w_{ij}\}$ be the extensions of $v_{i}$ to $L$ for $1\leq i\leq r$. Let $S$ be the normalization of $R$ in $L$ and $Y$ be the normalization of $X$ in $L$, with induced morphism $\varphi:Y\rightarrow X$.
Let $\{w_{ij}\}$ be the extensions of $v_{i}$ to $L$. Let $u_{i,j}$ be the canonical valuation of $\mathcal{O}_{w_{ij}}$, that is, a uniformizing parameter $t$ of $\mathcal{O}_{w_{i,j}}$ satisfies $u_{i,j}=1$. There exists a prime divisor $F_{ij}$ on $Y$ such that $\mathcal{O}_{w_{ij}}=\mathcal{O}_{F_{ij}}$. Write $w_{ij}=d_{ij}u_{ij}$. Let $D^{\prime}=\sum_{i,j}\frac{\lambda_{i}}{d_{ij}}F_{ij}$. Then
(21) $I(D^{\prime})_{n}=\cap_{i}\cap_{j}I(u_{ij})_{n\frac{\lambda_{i}}{d_{ij}}}\cap S=\cap_{i}\cap_{j}I(w_{ij})_{n\lambda_{i}}\cap S$
for all $n\in{\mathbb{N}}$.
3. A generalized McAdam theorem
In this section we prove a theorem (Theorem 3.5) which is the essential part of the proof of Theorem 1.5 stated in the introduction.
3.1. Resolutions and Alterations
The proof of [12, Section 5] shows the following.
Theorem 3.1.
([12, Section 5]) Let $R$ be a normal excellent local domain and let =$\mathcal{I}=\{I_{n}\}$ be a ${\mathbb{Q}}$-divisorial filtration on $R$. Suppose that there exists a projective birational morphism $\pi:X\rightarrow\mbox{Spec}(R)$ such that $X$ is nonsingular, and there exists an effective ${\mathbb{Q}}$-divisor $D$ on $X$ such that $\mathcal{I}=\mathcal{I}(D)$. If $m_{R}\in\mbox{Ass}(R/I_{m_{0}})$ for some $m_{0}\in{\mathbb{Z}}_{>0}$, then the analytic spread of $\mathcal{I}$ is $\ell(\mathcal{I})=\dim R$.
The following proposition is an extension of of [12, Section 6 ], with essentially the same proof.
Proposition 3.2.
([12, Section 6] Let $R\rightarrow S$ be a finite extension of reduced rings where $R$ is a Noetherian local ring of dimension $d$ and $S$ is a semi local ring. Let $\mathcal{J}=\{J_{n}\}$ be a graded filtration of $S$ and let =$\mathcal{I}=\{I_{n}\}$ be the graded filtration of $R$ defined by $I_{n}=J_{n}\cap R$. For $q$ a prime ideal of $S$, let $\mathcal{J}_{q}$ be the graded filtration $\mathcal{J}_{q}=\{(J_{n})_{q}\}$ of $S_{q}$. Suppose that $m_{R}\in\mbox{Ass}_{R}(R/I_{m_{0}})$ for some $m_{0}>0$. Then there exists a maximal ideal $q$ of $S$ such that $qS_{q}\in\mbox{Ass}_{S_{q}}(S_{q}/(J_{m_{0}})_{q})$. If we further have that $R$ and $S$ are equidimensional of dimension $d$, $\ell(\mathcal{J}_{q})=d$ and there exists a nonzero divisor of $R$ in $M=\mbox{Ann}_{R}(S/R)$, then $\ell(\mathcal{I})=d$.
Proof.
Let $\{\mathfrak{m}_{i}$} be the maximal ideals of $S$. Since we assume that $m_{R}\in\mbox{Ass}_{R}(R/I_{m_{0}})$, there exists $y\in R/I_{n_{0}}$ such that $m_{R}=\mbox{ann}_{R}(y)$. We have that $y\in S/J_{n}$ is nonzero since $R/I_{m_{0}}\rightarrow S/J_{m_{0}}$ is an inclusion. Thus $\mbox{Ann}_{S}(y)\neq S$. Since maximal elements in the set of annihilators of elements of $S/J_{m_{0}}$ are prime ideals (by [30, Theorem 6.1]), there exists a prime ideal $q$ in $S$ which contains $\mbox{Ann}_{S}(y)$ and is the annihilator of an element $z$ of $S/J_{m_{0}}$. We have that $q$ contains $m_{R}S$ and
(22) $\sqrt{m_{R}S}=\cap\mathfrak{m}_{i}$
so $q$ is a maximal ideal of $S$ such that $q\in\mbox{Ass}_{S}(S/J_{m_{0}})$, and thus $qS_{q}\in\mbox{Ass}_{S_{q}}(S_{q}/(J_{m_{0}})_{q})$.
Further suppose that
(23) $\ell(\mathcal{J}_{q})=\dim(S_{q})=\dim R$
and there exists a nonzero divisor of $R$ in $M=\mbox{Ann}_{R}(S/R)$
Let $B=\oplus_{n\geq 0}J_{n}$, which is a graded ring. We thus have by (2), and (23) that $\dim B/m_{R}B=\dim R$. Thus there is a chain of distinct prime ideals
$C_{0}\subset C_{1}\subset C_{2}\subset\cdots\subset C_{d}$
in $B$ which contain $m_{R}B$, where $d=\dim R$.
There is a natural inclusion of graded rings $R[\mathcal{I}]=\oplus_{n\geq 0}I_{n}\subset B=\oplus_{n\geq 0}J_{n}$. We will now show that $B$ is integral over $R[\mathcal{I}]$. It suffices to show that homogeneous elements of $B$ are integral over $R[\mathcal{I}]$.
For $a\in{\mathbb{Z}}_{>0}$, let $R[\mathcal{I}]_{a}$ be the $a$-th truncation of $R[\mathcal{I}]$ and $B_{a}$ be the $a$-th truncation of $B$, so that $R[\mathcal{I}]_{a}$ is the subalgebra of $R[\mathcal{I}]$ generated by $\oplus_{n\leq a}I_{n}$ and $B_{a}$ is the subalgebra of $B$ generated by $\oplus_{n\leq a}J_{n}$. Suppose that $f\in J_{a}$ for some $a$. Then $f\in B_{a}$. Let $0\neq x$ be a nonzero divisor of $R$ in $\mbox{Ann}_{R}(S/R)$. Then $xJ_{n}\subset I_{n}$ for all $n$ since $I_{n}=J_{n}\cap R$. Thus $xB_{a}\subset R[\mathcal{I}]_{a}$, so $f^{i}\in\frac{1}{x}R[\mathcal{I}]_{a}$ for all $i\in{\mathbb{N}}$, and so the algebra $R[\mathcal{I}]_{a}[f]\subset\frac{1}{x}R[\mathcal{I}]_{a}$. Since $\frac{1}{x}R[\mathcal{I}]_{a}$ is a finitely generated $R[\mathcal{I}]_{a}$-module and $R[\mathcal{I}]_{a}$ is a Noetherian ring, the ring $R[\mathcal{I}]_{a}[f]$ is a finitely generated $R[\mathcal{I}]_{a}$-module, so that $f$ is integral over $R[\mathcal{I}]_{a}$.
We have a chain of prime ideals
$Q_{0}\subset Q_{1}\subset Q_{2}\subset\cdots\subset Q_{d}$
in $R[\mathcal{I}]$ where $Q_{i}:=C_{i}\cap R[\mathcal{I}]$. The $Q_{i}$ are all distinct since the $C_{i}$ are all distinct and $B$ is integral over $R[\mathcal{I}]$ (by [3, Theorem A.6 (b)]). We have that $m_{R}R[\mathcal{I}]\subset Q_{0}$, so that $\dim R[\mathcal{I}]/m_{R}R[\mathcal{I}]\geq d$. Since this is the maximum possible dimension of $R[\mathcal{I}]/m_{R}R[\mathcal{I}]$ by (2), we have that $\ell(\mathcal{I})=d$.
Lemma 3.3.
Suppose that $R$ is an excellent local domain and $v_{1},\ldots,v_{r}$ are divisorial valuations in $R$. Then there exists a birational projective morphism $\pi:X\rightarrow\mbox{Spec}(R)$ such that $X$ is normal and there exist prime divisors $F_{1},\ldots,F_{r}$ on $X$ such that $\mathcal{O}_{X,F_{i}}$ is the valuation ring of $v_{i}$ for $1\leq i\leq r$.
Proof.
This follows from [13, Remark 6.6 to Lemma 6.5]. ∎
Definition 3.4.
([24, Definition 2.20]) Let $X$ be a Noetherian integral scheme. An Alteration $X^{\prime}$ of $X$ is an integral scheme $X^{\prime}$ with a morphism $\varphi:X^{\prime}\rightarrow X$ which is dominant, proper and such that for some open $U\subset X$, $\varphi^{-1}(U)\rightarrow U$ is finite.
A scheme $X$ is nonsingular if $\mathcal{O}_{X,q}$ is a regular local ring for all $q\in X$. An alteration $\varphi:X^{\prime}\rightarrow X$ is regular if $X^{\prime}$ is nonsingular.
3.2. Associated primes and maximal analytic spread
In this subsection we prove the following theorem. Theorem 3.5 is proven for excellent equicharacteristic zero local domains in Theorem 1.3 [12].
Theorem 3.5.
Let $R$ be an equidimensional excellent local ring which has the property that if $P$ is a minimal prime of $R$ and $X\rightarrow\mbox{Spec}(R/P)$ is a birational projective morphism, with $X$ integral, then there exists a projective regular alteration $X^{\prime}\rightarrow X$. Let =$\mathcal{I}=\{I_{n}\}$ be a ${\mathbb{Q}}$-divisorial filtration on $R$. If $m_{R}\in\mbox{Ass}(R/I_{m_{0}})$ for some $m_{0}\in{\mathbb{Z}}_{>0}$, then the analytic spread of $\mathcal{I}$ is $\ell(\mathcal{I})=\dim R$.
Proof.
Suppose that $m_{R}\in\mbox{Ass}_{R}(R/I_{m_{0}})$ for some $m_{0}$. Let $d=\dim R$.
Step 1. Suppose that $R$ is normal. By Lemma 3.3, there exists a birational projective morphism $Y_{1}\rightarrow\mbox{Spec}(R)$ such that $Y_{1}$ is normal and there exists an effective ${\mathbb{Q}}$-Weil divisor $D_{1}$ on $Y_{1}$ such that $\mathcal{I}=\mathcal{I}(D_{1})$. By the assumptions of the theorem, there exists a projective regular alteration $Z\rightarrow Y$.
Let $K$ be the quotient field of $R$ and let $L=\mathcal{O}_{Z,\eta}$ where $\eta$ is the generic point of $Z$. Let $S$ be the normalization of $R$ in $L$, with natural projective birational morphism $Z\rightarrow\mbox{Spec}(S)$. $S$ is a finitely generated $R$-module since $R$ is excellent. Let $q_{1},\ldots,q_{e}$ be the distinct maximal ideals of $S$. Let $\mathcal{I}(R)=\mathcal{I}=\mathcal{I}(D_{1})$. Let $Z_{1}\rightarrow Y_{1}$ be the normalization of $Y_{1}$ in $L$, with projective birational morphisms $Z\rightarrow Z_{1}\rightarrow\mbox{Spec}(S)$. Define $\mathcal{I}(S)$ from $\mathcal{I}(R)$ by the construction of (11). Define a ${\mathbb{Q}}$-Weil divisor $D_{1}^{\prime}$ on $Z_{1}$ from $D_{1}$ by the construction of (21), so that $\mathcal{I}(S)=\mathcal{I}(D_{1}^{\prime})$. Writing $D_{1}^{\prime}=\sum b_{ij}F_{ij}$, let $D^{\prime}$ be the ${\mathbb{Q}}$-divisor $D^{\prime}=\sum b_{ij}\overline{F}_{ij}$, where $\overline{F}_{ij}$ is the unique codimension 1 subvariety of $Z$ which dominates the codimension one subvariety $F_{ij}$ in the normal variety $Z_{1}$. Then $\mathcal{I}(D^{\prime})=\mathcal{I}(D_{1}^{\prime})=\mathcal{I}(S)$.
Since $m_{R}\in\mbox{Ass}(R/I(R)_{m_{0}})$, $I(R)_{m_{0}}:m_{R}^{\infty}\neq I(R)_{m_{0}}$ which implies that $I(S)_{m_{0}}:J(S)^{\infty}\neq I(S)_{m_{0}}$ by Lemma 2.3. Thus there exists some $q\in\{q_{1},\ldots q_{e}\}$ such that $[I(S)_{m_{0}}:J(S)^{\infty}]_{q}\neq[I(S)_{m_{0}}]_{q}$.
Define the divisorial filtration $\mathcal{I}(S_{q})=\{I(S_{q})_{n}\}$ on the $d$-dimensional excellent normal local ring $S_{q}$ by the localization $I(S_{q})_{n}=I(S)_{n}\otimes_{S}S_{q}$. We have that $I(S_{q})_{m_{0}}:m_{S_{q}}^{\infty}\neq I(S_{q})_{m_{0}}$. The base change $Z_{q}=Z\times_{S}S_{q}\rightarrow\mbox{Spec}(S_{q})$ is a birational projective morphism such that $Z_{q}$ is nonsingular. Further, $\mathcal{I}(S_{q})=\mathcal{I}(D^{\prime}|Z_{q})$ for the ${\mathbb{Q}}$-Cartier divisor $D^{\prime}|Z_{q}$ on $Z_{q}$. By Theorem 3.1, $\ell(\mathcal{I}(S_{q}))=\dim S=d$. $S[\mathcal{I}(S_{q})]=S[\mathcal{I}(S)]\otimes_{S}S_{q}$ and $\dim S_{q}[\mathcal{I}(S_{q})]/q_{q}S_{q}[\mathcal{I}(S_{q})]=\ell(\mathcal{I}(S_{q}))=d$ implies there exists a chain of distinct prime ideals $P_{0}\subset P_{1}\subset\cdots\subset P_{d}$ in $S[\mathcal{I}(S)]$ such that $q\subset P_{0}$. From the inclusion $R[\mathcal{I}(R)]\subset S[\mathcal{I}(S)]$, we have a chain of prime ideals $Q_{0}=P_{0}\cap R[\mathcal{I}(R)]\subset Q_{1}=P_{1}\cap R[\mathcal{I}(R)]\subset\cdots\subset Q_{d}=P_{d}\cap R[\mathcal{I}(R)]$. Now $R[\mathcal{I}(R)]\rightarrow S[\mathcal{I}(S)]$ is an integral extension by Proposition 2.4. Thus the prime ideals $Q_{0}\subset Q_{1}\subset\cdots\subset Q_{d}$ are all distinct by [3, Theorem A.6]. Since these ideals all contain $q\cap R=m_{R}$, we have that $\ell(\mathcal{I}(R))\geq d$. Since we have the upper bound $\ell(\mathcal{I}(R))\leq d$ of (2), we have that $\ell(\mathcal{I}(R))=d$. This completes the proof of Theorem 3.5 in the case that $R$ is normal.
Step 2. Now suppose that $R$ is a domain, which is not normal. There exist divisorial valuations $\nu_{1},\ldots,\nu_{t}$ of $R$ and $a_{1},\ldots,a_{t}\in{\mathbb{Q}}_{>0}$ such that =$\mathcal{I}=\{I_{n}\}$ where $I_{n}=I_{R}(\nu_{1})_{a_{1}n}\cap\cdots\cap I_{R}(\nu_{t})_{a_{t}n}$ for $n\geq 0$. Let $S$ be the normalization of $R$ in the quotient field of $R$. Let $J(\nu_{i})_{m}=\{f\in S\mid\nu_{i}(f)\geq m\}$. For $n\in{\mathbb{N}}$, let
$J_{n}=J(\nu_{1})_{a_{1}n}\cap\cdots\cap J(\nu_{t})_{a_{t}n}$
so that $J_{n}\cap R=I_{n}$. Let $\mathcal{J}=\{J_{n}\}$. By Proposition 3.2, there exists a maximal ideal $q$ of $S$ such that $qS_{q}\in\mbox{Ass}_{S_{q}}(S_{q}/(J_{m_{0}})_{q})$. Now $\mathcal{J}_{q}$ is a ${\mathbb{Q}}$-divisorial filtration of the $d$-dimensional normal local ring $S_{q}$. A birational projective morphism $X\rightarrow\mbox{Spec}(S)$ induces a birational projective morphism $X\rightarrow\mbox{Spec}(R)$, so $\ell(\mathcal{J}_{q})=d$ by Step 1. Let $0\neq x\in R$ be in the conductor of $S$ over $R$, so that $x\in\mbox{Ann}_{R}(S/R)$ is a nonzero divisor on $R$. Then by Proposition 3.2, we have that $\ell(\mathcal{I})=d$.
Step 3. Suppose that $R$ is reduced and equidimensional and $\mathcal{I}$ has an expression (16). Let $S=\oplus_{i=1}^{t}R/P_{i}$ where $P_{i}$, $1\leq i\leq t$ are the minimal primes of $R$. Let $\mathcal{J}=\{J_{n}\}$ where
$J_{n}=\oplus_{i=1}^{t}[\cap_{j}I_{R/P_{i}}(v_{ij})_{n\lambda_{ij}}].$
We have that $J_{n}\cap R=I_{n}$. By Proposition 3.2, there exists a maximal ideal $q$ of $S$ such that $qS_{q}\in\mbox{Ass}_{S_{q}}(S_{q}/(J_{m_{0}})_{q})$.
Let $M=S/R$, $A=\mbox{Ann}_{R}M$. Localizing the exact sequence of $R$-modules
$0\rightarrow R\rightarrow\oplus R/P_{i}\rightarrow M\rightarrow 0$
at $P_{i}$, we see that $M_{P_{i}}=0$ for all $i$. Thus $A_{P_{i}}=R_{P_{i}}$ for all $i$ so that $A\not\subset P_{i}$ for all $i$, and so $A\not\subset\cup P_{i}$ by prime avoidance. Thus there exists $x\in A\setminus\cup P_{i}$. Since $R$ is reduced, $x$ is a nonzero divisor of $R$. Now $S_{q}\cong(R/P_{i})_{m_{R/P_{i}}}$ for some $i$, and $\mathcal{J}_{q}$ is a ${\mathbb{Q}}$-divisorial filtration on the $d$-dimensional local domain $S_{q}$, so that $\ell(\mathcal{J}_{q})=d$ by Step 2. Thus $\ell(\mathcal{I})=d$ by Proposition 3.2 (this is where the assumption that $R$ is equidimensional is used).
Step 4. Now suppose that $R$ is an arbitrary equidimensional local ring. Let =$\mathcal{I}=\{I_{n}\}$ have an expression (16). Let $\tilde{R}=R/\cap_{i=1}^{t}P_{i}$, which is a reduced local ring, with natural surjection $R\rightarrow\tilde{R}$. Since $P_{i}$ is contained in the kernel of $\varphi_{i}$ for all $i$, we have that $\cap_{i=1}^{t}P_{i}$ is contained in $\varphi^{-1}(I_{R/P_{i}}(v_{ij})_{n\lambda_{i,j}})$ for all $i,j$, and so $\cap_{i=1}^{t}P_{i}$ is contained in $I_{n}$ for all $n$. Let $\alpha_{i}:\tilde{R}\rightarrow R/P_{i}$ be the natural projection for $1\leq i\leq t$. Let $\tilde{\mathcal{I}}$ be the ${\mathbb{Q}}$-divisorial filtration on $\tilde{R}$ defined by $\tilde{I}_{n}=\cap_{i=1}^{t}\cap_{i,j}\alpha_{i}^{-1}(I_{R/P_{i}}(v_{ij})_{n\lambda_{ij}})$. We have that $R/I_{n}\cong\tilde{R}/\tilde{I}_{n}$ for all $n$ and $m_{\tilde{R}}=m_{R}\tilde{R}$. Thus $m_{R}$ an associated prime of $R/I_{n}$ implies $m_{\tilde{R}}$ is an associated prime of $\tilde{R}/\tilde{I}_{n}$, so (~$\ell(\tilde{\mathcal{I}})=\dim\tilde{R}=\dim R$ by Step 3. Since there is a natural surjection
$R[\mathcal{I}]/m_{R}R[\mathcal{I}]\rightarrow\tilde{R}[\tilde{\mathcal{I}}]/m_{\tilde{R}}\tilde{R}[\tilde{\mathcal{I}}],$
we have that $\ell(\mathcal{I})\geq\dim R$, so by (2), we have that $\ell(\mathcal{I})=\dim R$ and the theorem holds for $R$.
4. A divisorial valuation whose first difference function doesn’t have a limit
4.1. Boundedness of the difference function
In this subsection we prove the following lemma.
Lemma 4.1.
Suppose that $R$ is a $d$-dimensional normal local ring which is essentially of finite type over a field and =$\mathcal{I}=\{I_{n}\}$ is a ${\mathbb{Q}}$-divisorial filtration of $m_{R}$-primary ideals on $R$. Then there exists a constant $c$ such that
$\lambda_{R}(I_{n}/I_{n+1})\leq cn^{d-1}$
for all $n$. Thus
$\limsup_{n\rightarrow\infty}\frac{\lambda_{R}(I_{n}/I_{n+1})}{n^{d-1}}$
is finite.
Proof.
There exists a field $\tilde{k}$ such that $\tilde{k}\subset R$ and $[R/m_{R}:\tilde{k}]<\infty.$ Let $\mathcal{I}(R)=\mathcal{I}$. The proof of Step 1 of Theorem 3.5 in Section 3.2 and [24, Theorem 4.1], shows that there exists a finite field extension $L$ of the quotient field $K$ of $R$ such that if $S$ is the normalization of $R$ in $L$ then there exists a resolution of singularities $Z\rightarrow\mbox{Spec}(S)$ and an effective ${\mathbb{Q}}$-divisor $D^{\prime}$ on $Z$ such that $\mathcal{I}(S)=\mathcal{I}(D^{\prime})$ and $I(S)_{n}\cap R=I(R)_{n}=I_{n}$ for all $n$. Thus we have inclusions
(24) $I_{n}/I_{n+1}\subset I(S)_{n}/I(S)_{n+1}$
for all $n$. There exists a positive integer $a$ such that $aD^{\prime}$ is a ${\mathbb{Z}}$-divisor. Write $n=m(n)a+b(n)$n where $0\leq b(n)<a$. Let $F_{n}=\lceil nD^{\prime}\rceil-m(n)aD^{\prime}$. The $\{F_{n}\}$ are effective divisors such that set of all $F_{n}$ is a finite set. Let
$G_{n}=\lceil(n+1)D^{\prime}\rceil-\lceil nD^{\prime}\rceil.$
Again, the set of all $G_{n}$ is finite. We have short exact sequences
$0\rightarrow\mathcal{O}_{Z}(-G_{n})\rightarrow\mathcal{O}_{Z}\rightarrow\mathcal{O}_{G_{n}}\rightarrow 0$
yeilding short exact sequences
$0\rightarrow\mathcal{O}_{Z}(-\lceil(n+1)D^{\prime}\rceil)\rightarrow\mathcal{O}_{Z}(-\lceil nD^{\prime}\rceil)\rightarrow\mathcal{O}_{G_{n}}\otimes\mathcal{O}_{Z}(-m(n)aD^{\prime}-F_{n})\rightarrow 0$
Taking global sections, we see that
(25) $\begin{array}[]{lll}I(S)_{n}/I(S)_{n+1}&=&\Gamma(Z,\mathcal{O}_{Z}(-\lceil(n+1)D^{\prime}\rceil))/\Gamma(Z,\mathcal{O}_{Z}(-\lceil nD^{\prime}\rceil))\\ &\subset&\Gamma(Z,\mathcal{O}_{G_{n}}\otimes_{\mathcal{O}_{Z}}\mathcal{O}_{Z}(-m(n)aD^{\prime}-F_{n})).\end{array}$
Let $A$ be an ample divisor on the projective $R$-scheme $Z$ such that $-D^{\prime}<A$. Let $F^{\prime}\in\{F_{n}\}$ and $G^{\prime}\in\{G_{n}\}$. $G^{\prime}$ is a projective $\tilde{k}$-scheme of dimension $d-1$. For a coherent sheaf $\mathcal{F}$ on $G^{\prime}$ let $h^{0}(\mathcal{F})=\dim_{\tilde{k}}\Gamma(G^{\prime},\mathcal{F})$. Then
$h^{0}(\mathcal{F})=[R/m_{R}:\tilde{k}]\ell_{R}(\Gamma(G^{\prime},\mathcal{F})).$
We have that
$h^{0}(\mathcal{O}_{G^{\prime}}\otimes\mathcal{O}_{Z}(-laD-F^{\prime}))\leq h^{0}(\mathcal{O}_{G^{\prime}}\otimes\mathcal{O}_{Z}(lA-F^{\prime}))=\chi(\mathcal{O}_{G^{\prime}}\otimes\mathcal{O}_{Z}(lA-F^{\prime}))$
for $l\gg 0$ by Proposition III.5.3 [22] or Theorem 17.8 [9]. The Euler characteristic $\chi(\mathcal{O}_{G^{\prime}}\otimes\mathcal{O}_{Z}(lA-F^{\prime}))$ is a polynomial in $l$ of degree $\leq\dim G^{\prime}=d-1$ by [38], [26] or Theorem 19.1 [9]. Thus there exists a constant $c^{\prime}$ such that
(26) $h^{0}(\mathcal{O}_{G_{n}}\otimes\mathcal{O}_{Z}(-m(n)aD-F_{n}))\leq c^{\prime}n^{d-1}$
for all positive $n$, and so, there exists a positive constant $c$ such that
$\lambda_{R}(I_{n}/I_{n+1})\leq cn^{d-1}$
for all $n$ by (24), (25) and (26).
4.2. Beatty Sequences
For $x\in{\mathbb{R}}$, $\lfloor x\rfloor$ will denote the round down of $x$ and $\lceil x\rceil$ will denote the round up of $x$. The fractional part of $x$ is $\{x\}=x-\lfloor x\rfloor$.
Let $\alpha\in{\mathbb{R}}$. The Beatty sequence of $\alpha$ is the sequence $\sigma_{\alpha}(n)=\lfloor\alpha(n+1)\rfloor-\lfloor\alpha n\rfloor$ for $n\in{\mathbb{Z}}_{>0}$. We remark that if $\alpha$ is irrational, then $\sigma_{\alpha}(n)=\lceil\alpha(n+1)\rceil-\lceil\alpha n\rceil$ for all $n\in{\mathbb{Z}}_{>0}$. We have the following classical result.
Proposition 4.2.
Suppose that $\alpha$ is an irrational number. Then there are exactly two values of $\sigma_{\alpha}(n)$ for $n\in{\mathbb{Z}}_{>0}$; they are $\lfloor\alpha\rfloor$ and $\lceil\alpha\rceil$. Each of these two values of $\sigma_{\alpha}(n)$ occur for infinitely many values of $n\in{\mathbb{Z}}_{>0}$.
We will give a short proof of Proposition 4.2 in this subsection.
Definition 4.3.
Let $x_{n}\in{\mathbb{R}}$ be a sequence. The sequence $x_{n}$ is uniformly distributed Mod 1 if for every $a,b$ with $0\leq a<b\leq 1$, we have that
$\lim_{n\rightarrow\infty}\frac{1}{n}{\rm card}\{0\leq j\leq n-1\mid\{x_{j}\}\in[a,b]\}=b-a.$
A proof of the following theorem is given in formula (1.1) of page 9 of [35].
Theorem 4.4.
Let $\alpha\in{\mathbb{R}}$ be irrational. Then the sequence $n\alpha$ for $n\in{\mathbb{N}}$ is uniformly distributed mod 1.
We now give a proof of Proposition 4.2, following the note [5].
We have that $\alpha(n+1)-1<\lfloor\alpha(n+1)\rfloor<\alpha(n+1)$ and $\alpha n-1<\lfloor\alpha n\rfloor<\alpha n$. Thus
$\alpha(n+1)-1-\alpha n<\lfloor\alpha(n+1)\rfloor-\lfloor\alpha n\rfloor<\alpha(n+1)-(\alpha n-1)$
and so
$\alpha-1<\lfloor\alpha(n+1)\rfloor-\lfloor\alpha n\rfloor<\alpha+1.$
Since $\lfloor\alpha(n+1)\rfloor-\lfloor\alpha n\rfloor$ is an integer and $\alpha$ is not, $\lfloor\alpha(n+1)\rfloor-\lfloor\alpha n\rfloor$ is equal to $\lfloor\alpha\rfloor$ or $\lceil\alpha\rceil$. We have that
(27) $\lfloor\alpha(n+1)\rfloor-\lfloor\alpha n\rfloor=\alpha(n+1)-\{\alpha(n+1)\}-(\alpha n-\{\alpha n\})=\alpha+\{\alpha n\}-\{\alpha n+\alpha\}.$
For $x,y\in{\mathbb{R}}$, we have that
$\{x+y\}=\left\{\begin{array}[]{l}\{x\}+\{y\}\mbox{ if }\{x\}+\{y\}<1\\ \{x\}+\{y\}-1\mbox{ if }\{x\}+\{y\}\geq 1.\end{array}\right.$
By Theorem 4.4, choosing $a,b$ such that $11-\{\alpha\}<a<b<$ 1, we find that there are infinitely many $n\in{\mathbb{N}}$ such that $\{\alpha n\}+\{\alpha\}>1$. Hence,
$\begin{array}[]{lll}\sigma_{\alpha}(n)&=&\lfloor\alpha(n+1)\rfloor-\lfloor\alpha n\rfloor=\alpha+\{\alpha n\}-\{\alpha n+\alpha\}=\alpha+\{\alpha n\}-(\{\alpha n\}+\{\alpha\}-1)\\ &=&\alpha-\{\alpha\}+1=\lceil\alpha\rceil.\end{array}$
Hence $\lceil\alpha\rceil$ appears infinitely many times in the sequence $\sigma_{\alpha}(n)$.
Again by Theorem 4.4, choosing $a,b$ so that $0<a<b<1-\{\alpha\}$, we find there are infinitely many $n\in{\mathbb{N}}$ such that $\{\alpha n\}+\{\alpha\}<1$. Continuing as in the above paragraph, we find that $\lfloor\alpha\rfloor$ occurs infinitely many times in the sequence $\sigma_{\alpha}(n)$.
4.3. Geometric preliminaries
Let $V$ be a projective variety over a field $k$ and $\mathcal{F}$ be a coherent sheaf on $V$. We will write $h^{i}(\mathcal{F})=\dim_{k}H^{i}(V,\mathcal{F})$. The Euler characteristic of $\mathcal{F}$ is
$\chi(\mathcal{F})=\sum_{i=0}^{\dim\mathcal{F}}(-1)^{i}h^{i}(\mathcal{F}).$
Similar statements to the following Lemma 4.5 are well known. For the readers convenience, we give a proof.
Lemma 4.5.
Suppose that $T$ is a nonsingular projective surface over a field of characteristic zero and $\{D_{n}\}_{n\in{\mathbb{Z}}_{>0}}$ is a family of ample divisors on $T$, having the property that if $C$ is a curve on $T$, then there exists a constant $\gamma(C)$ such that $(C\cdot D_{n})\leq\gamma(C)n$ for all $n\in{\mathbb{Z}}_{>0}$. Suppose that $G$ is a divisor on $T$. Then there exists a constant $c$ such that
$h^{1}(\mathcal{O}_{T}(D_{n}+G))\leq cn\mbox{ and }h^{2}(\mathcal{O}_{T}(D_{n}+G))\leq c\mbox{ for all $n\in{\mathbb{Z}}_{>0}$.}$
Proof.
By Kodaira vanishing,
(28) $H^{i}(T,\mathcal{O}_{T}(D_{n}+K_{T}))=0\mbox{ for all $n>0$ and $i>0$}$
where $K_{T}$ is a canonical divisor on $T$.
Suppose that $Z$ is an integral nonsingular curve on $T$. From the short exact sequence $0\rightarrow\mathcal{O}_{T}(-Z)\rightarrow\mathcal{O}_{T}\rightarrow\mathcal{O}_{Z}\rightarrow 0$, we have exact sequences
$\begin{array}[]{l}H^{1}(T,\mathcal{O}_{T}(D_{n}+K_{T}))\rightarrow H^{1}(T,\mathcal{O}_{T}(D_{n}+K_{T}+Z))\rightarrow H^{1}(Z,\mathcal{O}_{T}(D_{n}+K_{T}+Z)\otimes\mathcal{O}_{Z}))\\ \rightarrow H^{2}(T,\mathcal{O}_{T}(D_{n}+K_{T}))\rightarrow H^{2}(T,\mathcal{O}_{T}(D_{n}+K_{T}+Z))\rightarrow 0.\end{array}$
Since $((D_{n}+K_{T}+Z)\cdot Z)$ is bounded from below (as the $D_{n}$ are ample) we have that $h^{1}(\mathcal{O}_{T}(D_{n}+K_{T}+Z)\otimes\mathcal{O}_{Z}))$ is bounded by Lemma 3.1 [11]. Thus using equation (28), there exists a constant $c$ such that
(29) $h^{1}(\mathcal{O}_{T}(D_{n}+K_{T}+Z))\leq c\mbox{ and }h^{2}(\mathcal{O}_{T}(D_{n}+K_{T}+Z))=0\mbox{ for all }n>0.$
Now suppose that $G^{\prime}$ is a divisor on $T$. By Bertini’s theorem, we have a linear equivalence $G^{\prime}\sim A-B$ where $A$ and $B$ are ample effective divisors and $A$ and $B$ are integral nonsingular curves on $S$. From the short exact sequence
$0\rightarrow\mathcal{O}_{T}(-B)\rightarrow\mathcal{O}_{T}\rightarrow\mathcal{O}_{B}\rightarrow 0,$
we have exact sequences
$\begin{array}[]{l}H^{0}(B,\mathcal{O}_{T}(D_{n}+K_{T}+A)\otimes\mathcal{O}_{B})\rightarrow H^{1}(T,\mathcal{O}_{T}(D_{n}+K_{T}+G^{\prime}))\\ \rightarrow H^{1}(T,\mathcal{O}_{T}(D_{n}+K_{T}+A))\rightarrow H^{1}(B,\mathcal{O}_{T}(D_{n}+K_{T}+A)\otimes\mathcal{O}_{B})\\ \rightarrow H^{2}(T,\mathcal{O}_{T}(D_{n}+K_{T}+G^{\prime}))\rightarrow H^{2}(T,\mathcal{O}_{T}(D_{n}+K_{T}+A))\rightarrow 0.\end{array}$
By the Riemann Roch theorem on a nonsingular curve (formula (6) and Lemma 3.1 [10], Section 7.3 [29] or Section IV.1 of [22] over an algebraically closed field), Lemma 3.1 [11] and our assumption that $\deg(\mathcal{O}_{T}(D_{n})\otimes\mathcal{O}_{B})\leq\gamma(B)n$, we have that there exists a constant $c^{\prime}$ such that $h^{0}(B,\mathcal{O}_{T}(D_{n}+K_{T}+A)\otimes\mathcal{O}_{B})\leq c^{\prime}n$ for all $n$. Now using formula (29), we see that there is a constant $c_{1}$ such that
(30) $h^{1}(\mathcal{O}_{T}(D_{n}+K_{T}+G^{\prime})\leq c_{1}n\mbox{ and }h^{2}(\mathcal{O}_{T}(D_{n}+K_{T}+G^{\prime}))\leq c_{1}$
for all $n\in{\mathbb{N}}$.
Let $G$ be a divisor on $T$. Take $G^{\prime}=G-K_{T}$ in (30), and the conclusions of the lemma follow. ∎
4.4. The construction
We recall the construction of Section 4 of [10]. A commutative diagram of projective varieties
$\begin{array}[]{rll}Y\\ \tau\downarrow&\searrow\\ X&\stackrel{{\scriptstyle\lambda}}{{\rightarrow}}&\overline{Z}\\ \pi\downarrow\\ S\end{array}$
is constructed, where $X$ is a 3-fold which is a projective bundle over an Abelian surface $S$ (which we assume is over an algebraically closed field $k$ of characteristic zero) and $\tau$ is the blowup of a nonsingular curve $C$ lying on the zero section $S_{0}$ of $\pi$, with exceptional divisor $F$. $\tau$ induces a morphism $\overline{\tau}:F\rightarrow C$, realizing $F$ as a ruled surface over $C$. Let $\overline{S}$ be the strict transform of $S_{0}$, which is isomorphic to $S$. Let $\overline{Z}$ be the normal projective 3-fold which is obtained by contracting $S_{0}$ (it is shown in [10] that there is a morphism $X\rightarrow\overline{Z}$ with this property). Let $q$ be the closed (singular) point of $\overline{Z}$ which is the image of $S_{0}$. In [10] it is shown that $\lambda:X\rightarrow\overline{Z}$ is the blowup of $\lambda_{*}\mathcal{O}_{X}(-mS_{0})$ for some large $m$, which is an ideal sheaf whose cosupport is $\lambda(S_{0})=q$. Since $\tau:Y\rightarrow X$ is the blowup of an ideal sheaf whose cosupport $C$ is contained in $S_{0}$, we have that $Y\rightarrow\overline{Z}$ is the blowup of an ideal sheaf whose cosupport is $q$.
Let $R=\mathcal{O}_{\overline{Z},q}$ and let $U=Y\times_{\overline{Z}}\mbox{Spec}(R)$ with induced projective morphism $U\rightarrow\mbox{Spec}(R)$. By the above, $U$ is the blowup of an $m_{R}$-primary ideal $I$ of $R$.
4.5. The example
By the formula before Lemma 2.2 [10] and Theorem 4.1 [10], we have that
$I(nF):=\Gamma(U,\mathcal{O}_{U}(-nF))=\Gamma(U,\mathcal{O}_{U}(-\lceil n\frac{3}{9-\sqrt{3}}\rceil\overline{S}-nF)).$
Let $D_{n}=\lceil n\frac{3}{9-\sqrt{3}}\rceil\overline{S}+nF$ for $n\in{\mathbb{Z}}_{>0}$. The closure of the ample cone of $\overline{S}$ is equal to the closure of the nef cone of $\overline{S}$; that is, $\overline{\rm Amp}(\overline{S})=\overline{\rm Nef}(\overline{S})$. It is shown before formula (14) on page 12 of [10] that
(31) $\mathcal{O}_{U}(-D_{n})\otimes\mathcal{O}_{\overline{S}}\mbox{ is ample}.$
Let $C_{0}$ be the zero section of $\overline{\tau}:F\rightarrow C$. The closure of the effective cone of $F$ is $\overline{\rm Eff}(F)={\mathbb{R}}_{\geq 0}C_{0}+{\mathbb{R}}_{\geq 0}f$ by a formula on page 15 [10], where $f$ is the fiber by $\overline{\tau}$ of a closed point of $C$. It is shown on page 14 [10] that the class of $\mathcal{O}_{U}(F)\otimes\mathcal{O}_{F}$ in $({\rm Pic}(F)/\equiv)\otimes{\mathbb{R}}$ is that of $\mathcal{O}_{U}(-C_{0}-108f)$. Thus the class of $\mathcal{O}_{U}(-D_{n})\otimes\mathcal{O}_{F}$ in $({\rm Pic}(F)/\equiv)\otimes{\mathbb{R}}$ is that of $\mathcal{O}_{U}((n-\lceil n\frac{3}{9-\sqrt{3}})C_{0}+n108f)$.
On page 14 [10], the intersection products
$(C_{0}^{2})_{F}=-162,(C_{0}\cdot f)_{F}=1\mbox{ and }(f^{2})_{F}=0$
are computed. We may thus compute that
$((\mathcal{O}_{U}(-D_{n})\otimes\mathcal{O}_{F})\cdot C_{0})>0\mbox{ and }(\mathcal{O}_{U}(-D_{n})\otimes\mathcal{O}_{F})\cdot f)>0$
for $n\gg 0$, so that $((\mathcal{O}_{U}(-D_{n})\otimes\mathcal{O}_{F})^{2})>0$. Thus there exists $n_{0}$ such that
(32) $\mathcal{O}_{U}(-D_{n})\otimes\mathcal{O}_{F}\mbox{ is ample }$
for $n\geq n_{0}$ by the Nakai-Moishezon criterion (Theorem V.1.10 [22]).
Lemma 4.6.
There exists a constant $c^{\prime}$ such that
$\lambda_{R}(H^{1}(U,\mathcal{O}_{U}(-D_{n}))\leq c^{\prime}n\mbox{ and }\lambda_{R}(H^{2}(U,\mathcal{O}_{U}(-D_{n}))\leq c^{\prime}\mbox{ for all $n\geq 0$}.$
Proof.
As we observed at the beginning of this section, $U$ is the blowup of an $m_{R}$-primary ideal $I$, so there exists an effective divisor $A=a\overline{S}+bF$ such that $\mathcal{O}_{U}(-A)=I\mathcal{O}_{U}$ and so $-A$ is ample on $U$. There exists a sequence of surfaces $T_{1},T_{2},\ldots,T_{r}$ where $r=a+b$ such that each $T_{i}$ is one of $\overline{S}$ or $F$ and $A=T_{1}+\cdots+T_{r}$. We have short exact sequences
(33) $\begin{array}[]{l}0\rightarrow O_{U}(-T_{1})\otimes\mathcal{O}_{T_{2}}\rightarrow\mathcal{O}_{T_{1}+T_{2}}\rightarrow\mathcal{O}_{T_{1}}\rightarrow 0\\ \,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\vdots\\ 0\rightarrow\mathcal{O}_{U}(-(T_{1}+\cdots+T_{r-1}))\otimes\mathcal{O}_{T_{r}}\rightarrow\mathcal{O}_{A}\rightarrow\mathcal{O}_{T_{1}+\cdots+T_{r-1}}\rightarrow 0\\ 0\rightarrow\mathcal{O}_{U}(-A)\otimes\mathcal{O}_{T_{1}}\rightarrow\mathcal{O}_{A+T_{1}}\rightarrow\mathcal{O}_{A}\rightarrow 0\\ 0\rightarrow\mathcal{O}_{U}(-A-T_{1})\otimes\mathcal{O}_{T_{2}}\rightarrow\mathcal{O}_{A+T_{1}+T_{2}}\rightarrow\mathcal{O}_{A+T_{1}}\rightarrow 0\\ \,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\vdots\\ 0\rightarrow\mathcal{O}_{U}(-A-(T_{1}+\cdots+T_{r-1}))\otimes\mathcal{O}_{T_{r}}\rightarrow\mathcal{O}_{2A}\rightarrow\mathcal{O}_{A+T_{1}+\cdots+T_{r-1}}\rightarrow 0\\ \,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\vdots\\ \end{array}$
There exists $t_{0}$ such that $t\geq t_{0}$ implies $\mathcal{O}_{U}(-tA-(T_{1}+\cdots+T_{i-1}))\otimes\mathcal{O}_{T_{i}}(-K_{T_{i}})$ is ample on $T_{i}$, for $1\leq i\leq r$, where $K_{T_{i}}$ is a canonical divisor on $T_{i}$. Since $\mathcal{O}_{U}(-D_{n})\otimes\mathcal{O}_{T_{i}}$ are ample for all $i$ and $n\geq n_{0}$ by (31) and (32),
$H^{i}(T_{i},\mathcal{O}_{T}(-D_{n}-tA-(T_{1}+\cdots+T_{i-1}))\otimes\mathcal{O}_{T_{i}})=0$
for all $n\geq n_{0}$, $i$ and $t\geq t_{0}$ by Kodaira vanishing. From the exact sequences of (33), we see that
(34) $H^{i}(\mathcal{O}_{U}(-D_{n})\otimes\mathcal{O}_{mA+T_{1}+\cdots+T_{i}})\cong H^{i}(U,\mathcal{O}_{U}(-D_{n})\otimes\mathcal{O}_{t_{0}A})$
for $n\geq n_{0}$, $m\geq t_{0}$ and $11\leq i\leq r-$ 1. We may now replace $A$ with $t_{0}A$ and replace $r$ with $u=t_{0}r$ and consider the short exact sequences
(35) $\begin{array}[]{l}0\rightarrow\mathcal{O}_{T}(-T_{1})\otimes\mathcal{O}_{T_{2}}\rightarrow\mathcal{O}_{T_{1}+T_{2}}\rightarrow\mathcal{O}_{T_{1}}\rightarrow 0\\ \,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\vdots\\ 0\rightarrow\mathcal{O}_{T}(-(T_{1}+\cdots+T_{u-1}))\otimes\mathcal{O}_{T_{u}}\rightarrow\mathcal{O}_{A}\rightarrow\mathcal{O}_{T_{1}+\cdots+T_{u-1}}\rightarrow 0.\end{array}$
By (31), (32) and Lemma 4.5, there exists a positive constant $c_{2}$ such that
$h^{1}(\mathcal{O}_{T}(-D_{n})\otimes\mathcal{O}_{T_{1}})\leq c_{2}n,\,\,h^{1}(\mathcal{O}_{U}(-D_{n}-(T_{1}+\cdots+T_{i-1}))\otimes\mathcal{O}_{T_{i}}))\leq c_{2}n,$
$h^{2}(\mathcal{O}_{T}(-D_{n})\otimes\mathcal{O}_{T_{1}})\leq c_{2},\,\,h^{2}(\mathcal{O}_{T}(-D_{n}-(T_{1}+\cdots+T_{i-1}))\otimes\mathcal{O}_{T_{i}}))\leq c_{2}$
for $2\leq i\leq u$ and all $n\geq 0$.
Thus for $n\geq n_{0}$,
$\begin{array}[]{l}\lambda_{R}(\lim_{\leftarrow}H^{1}(U,\mathcal{O}_{U}(-D_{n})\otimes\mathcal{O}_{U}/\mathcal{O}_{U}(-mA)))\\ =\dim_{k}(\lim_{\leftarrow}H^{1}(U,\mathcal{O}_{U}(-D_{n})\otimes\mathcal{O}_{U}/\mathcal{O}_{U}(-mA)))\\ \leq uc_{2}n\end{array}$
and
$\begin{array}[]{l}\lambda_{R}(\lim_{\leftarrow}H^{2}(U,\mathcal{O}_{U}(-D_{n})\otimes\mathcal{O}_{U}/\mathcal{O}_{U}(-mA)))\\ =\dim_{k}(\lim_{\leftarrow}H^{2}(U,\mathcal{O}_{U}(-D_{n})\otimes\mathcal{O}_{U}/\mathcal{O}_{U}(-mA)))\\ \leq uc_{2},\end{array}$
where the inverse limits are over $m\in{\mathbb{Z}}_{>0}$. Now $I$ is $m_{R}$-primary and $I^{m}\mathcal{O}_{U}=\mathcal{O}_{U}(-mA)$, so that
$\lim_{\leftarrow}H^{i}(U,\mathcal{O}_{U}(-D_{n})\otimes_{R}R/m_{R}^{m})\cong\lim_{\leftarrow}H^{i}(U,\mathcal{O}_{U}(-D_{n})\otimes_{\mathcal{O}_{U}}\mathcal{O}_{U}/\mathcal{O}_{U}(-mA))$
where the limits are over $m\in{\mathbb{Z}}_{>0}$. Since the $R$-modules $H^{i}(U,\mathcal{O}_{U}(-D_{n}))$ are supported on $m_{R}$, The $m_{R}$-adic completions of these modules are isomorphic to themselves. By the theorem on formal functions, Theorem III.1.1 [22],
$H^{i}(U,\mathcal{O}_{U}(-D_{n}))\cong\lim_{\leftarrow}H^{i}(U,\mathcal{O}_{U}(-D_{n})\otimes_{R}R/m_{R}^{n})$
for $n\geq n_{0}$ and $i=1$ and 2. Now choose $c_{3}$ so that $\lambda_{R}H^{i}(U,\mathcal{O}_{U}(-D_{n}))\leq c_{3}$ for $n\geq n_{0}$ and $i=1,2$. Let $c^{\prime}=uc_{2}+c_{2}$, so the conclusions of the lemma hold.
We now establish that
(36) $\lambda_{R}(R/I(nF))=\chi(\mathcal{O}_{D_{n}})+\tau(n)\mbox{ for $n\in{\mathbb{N}}$}$
where $\tau(n)$ is a function such that there exists a constant $c$ such that $|\tau(n)|\leq cn$ for all $n\in{\mathbb{N}}$. We start with the short exact sequences
$0\rightarrow\mathcal{O}_{U}(-D_{n})\rightarrow\mathcal{O}_{U}\rightarrow\mathcal{O}_{D_{n}}\rightarrow 0,$
yielding, since the dimension of the preimage of $m_{R}$ in $U$ is two and $R$ is normal, exact sequences
$\begin{array}[]{l}0\rightarrow R/I(nF)\rightarrow H^{0}(D_{n},\mathcal{O}_{D_{n}})\rightarrow H^{1}(U,\mathcal{O}_{U}(-D_{n}))\rightarrow H^{1}(U,\mathcal{O}_{U})\rightarrow H^{1}(D_{n},\mathcal{O}_{D_{n}})\\ \rightarrow H^{2}(U,\mathcal{O}_{U}(-D_{n}))\rightarrow H^{2}(U,\mathcal{O}_{U})\rightarrow H^{2}(U,\mathcal{O}_{D_{n}})\rightarrow 0\end{array}$
giving the formula
$\begin{array}[]{lll}\lambda_{R}(R/I(nF))&=&\chi(\mathcal{O}_{D_{n}})-\lambda_{R}H^{1}(U,\mathcal{O}_{U}(-D_{n}))+\lambda_{R}H^{1}(U,\mathcal{O}_{U})\\ &&+\lambda_{R}H^{2}(U,\mathcal{O}_{U}(-D_{n}))-\lambda_{R}H^{2}(U,\mathcal{O}_{U}).\end{array}$
Formula (36) is now obtained from Lemma 4.6.
We can compute $\chi(\mathcal{O}_{D_{n}})$ by regarding $D_{n}$ as a closed subscheme of the projective variety $Y$. From the exact sequence
$0\rightarrow\mathcal{O}_{Y}(-D_{n})\rightarrow\mathcal{O}_{Y}\rightarrow\mathcal{O}_{D_{n}}\rightarrow 0,$
we obtain that
(37) $\chi(\mathcal{O}_{D_{n}})=\chi(\mathcal{O}_{Y})-\chi(\mathcal{O}_{Y}(-D_{n})).$
Now from the Riemann Roch theorem for three dimensional nonsingular varieties (Exercise A.6.7 [22])
(38) $\begin{array}[]{lll}\chi(\mathcal{O}_{Y}(-D_{n}))&=&\frac{1}{12}(-D_{n}\cdot(-D_{n}-K_{Y})\cdot(2(-D_{n})-K_{Y}))\\ &&+\frac{1}{12}((-D_{n})\cdot c_{2}(Y))+1-p_{a}(Y)\\ &=&-\frac{1}{6}(D_{n}^{3})-\frac{1}{4}(D_{n}^{2}\cdot K_{Y})-\frac{1}{12}(D_{n}\cdot K_{Y}^{2}+c_{2}(Y))+1-p_{a}(Y).\end{array}$
Thus from (36), we obtain that there is a function $\delta(n)$ such that
(39) $\lambda_{R}(R/I(nF))=\frac{1}{6}(D_{n}^{3})+\frac{1}{4}(D_{n}^{2}\cdot K_{Y})+\delta(n)$
where there is a constant $c_{1}$ such that $|\delta(n)|<c_{1}n$ for all $n\in{\mathbb{N}}$.
Let $\alpha=\frac{3}{9-\sqrt{3}}$, so that
$D_{n}=\lceil\alpha n\rceil\overline{S}+nF.$
We have
$\alpha=\frac{3}{9-\sqrt{3}}=\frac{9}{26}+\frac{\sqrt{3}}{26},\,\,\alpha^{2}=\frac{84}{26^{2}}+\frac{18}{26^{2}}\sqrt{3},\,\,\alpha^{3}=\frac{810}{26^{3}}+\frac{246}{26^{3}}\sqrt{3}.$
On page 16 of [10], it is shown that
(40) $(\overline{S}^{3})=468,\,\,(\overline{S}^{2}\cdot F)=-162,\,\,(\overline{S}\cdot F^{2})=54,\,\,(F^{3})=54.$
We compute that
(41) $\begin{array}[]{lll}(D_{n}^{3})&=&\lceil\alpha n\rceil^{3}(\overline{S}^{3})+3n\lceil\alpha n\rceil^{2}(\overline{S}^{2}\cdot F)+3n^{2}\lceil\alpha n\rceil(\overline{S}\cdot F^{2})+n^{3}(F^{3})\\ &=&468\lceil\alpha n\rceil^{3}-486n\lceil\alpha n\rceil^{2}+162n^{2}\lceil\alpha n\rceil+54n^{3}\end{array}$
and
(42) $(D_{n}^{2}\cdot K_{Y})=\lceil\alpha n\rceil^{2}(\overline{S}^{2}\cdot K_{Y})+2n\lceil n\alpha\rceil(\overline{S}\cdot F\cdot K_{Y})+n^{2}(F^{2}\cdot K_{Y}).$
Thus
(43) $\lim_{n\rightarrow\infty}\frac{(D_{n})^{3}}{n^{3}}=468\alpha^{3}-486\alpha^{2}+162\alpha+54=\frac{12042}{169}-\frac{27}{169}\sqrt{3}.$
We obtain
(44) $e(\mathcal{I}(F))=3!\left(\lim_{n\rightarrow\infty}\frac{\lambda_{R}(R/I(nF))}{n^{3}}\right)=3!(\alpha^{3}468-486\alpha^{2}+162\alpha+54)=\frac{72252}{169}-\frac{162}{169}\sqrt{3}.$
We give the numeric expression of 42, which is computed using the formulas (40) and other formulas from [10]. This formula is not necessary to compute the asymptotic behavior of the first difference function $\lambda_{R}(I(nF))/I((n+1)F)$.
(45) $(D_{n}^{2}\cdot K_{Y})=-792\lceil n\alpha\rceil^{2}+564n\lceil n\alpha\rceil-175n^{2}.$
We now calculate the first difference function.
(46) $\lambda_{R}(I(nF)/I((n+1)F))=\lambda_{R}(R/I((n+1)F))-\lambda_{R}(R/I(nF))=\frac{1}{6}\left((D_{n+1}^{3})-(D_{n}^{3})\right)+\varepsilon(n)$
where there is a constant $c^{\prime\prime}$ such that $|\varepsilon(n)|<c^{\prime\prime}n$ for $n\in{\mathbb{N}}$. We have
(47) $\begin{array}[]{lll}\left((D_{n+1}^{3})-(D_{n}^{3})\right)&=&468\left(\lceil\alpha(n+1)\rceil^{3}-\lceil\alpha n\rceil^{3}\right)-486\left((n+1)\lceil\alpha(n+1)\rceil^{2}-n\lceil\alpha n\rceil^{2}\right)\\ &&+162\left((n+1)^{2}\lceil\alpha(n+1)\rceil-n^{2}\lceil\alpha n\rceil\right)+54\left((n+1)^{3}-n^{3}\right)\\ &=&468(\sigma_{\alpha}(n)(\lceil\alpha(n+1)\rceil^{2}+\lceil\alpha(n+1)\rceil\lceil\alpha n\rceil+\lceil\alpha n\rceil^{2})\\ &&-486\left(\sigma_{\alpha}(n)(\lceil\alpha(n+1)\rceil n+\lceil\alpha n\rceil n)+\lceil\alpha(n+1)\rceil^{2}\right)\\ &&+162\left(\sigma_{\alpha}(n)n^{2}+\lceil\alpha(n+1)\rceil(2n+1)\right)+54(3n^{2}+3n+1)\end{array}$
We have that
$\lceil\alpha\rceil=\lceil\frac{3}{9-\sqrt{3}}\rceil=1\mbox{ and }\lfloor\alpha\rfloor=0.$
By Proposition 4.2, the positive integers are the disjoint union ${\mathbb{Z}}_{>0}=\Sigma_{1}\coprod\Sigma_{2}$ of two infinite sets, where $\sigma_{\alpha}(n)=0$ if $n\in\Sigma_{1}$ and $\sigma_{\alpha}(n)=1$ if $n\in\Sigma_{2}$.
The limit as $n\rightarrow\infty$ restricted to $n\in\Sigma_{1}$ of $\frac{\lambda_{R}(I(nF)/I((n+1)F)}{n^{2}}$ is thus
$\frac{1}{6}\left(-486\alpha^{2}+324\alpha+162\right)=\frac{144504}{6(26)^{2}}-\frac{324}{6(26)^{2}}\sqrt{3}$
and the limit as $n\rightarrow\infty$ restricted to $n\in\Sigma_{2}$ of $\frac{\lambda_{R}(I(nF)/I((n+1)F)}{n^{2}}$ is
$\frac{1}{6}\left(918\alpha^{2}-810\alpha+324\right)=\frac{106596}{6(26)^{2}}-\frac{4536}{6(26)^{2}}\sqrt{3}.$
These two limits, over $n$ in $\Sigma_{1}$ and $\Sigma_{2}$ respectively are not equal, so the limit
$\lim_{n\rightarrow\infty}\frac{\lambda_{R}(I(nF)/I((n+1)F)}{n^{2}}$
does not exist.
5. Symbolic powers with infinitely many Rees valuations
In this section we construct an example of a height one prime ideal in a two dimensional normal local ring $R$ such that for all $n\geq 1$, the symbolic power $P^{(n)}$ has a Rees valuation other then the $PR_{P}$-adic valuation $v_{P}$, but for $1\leq m<n$, $v_{P}$ is the only common Rees valuation of $P^{(m)}$ and $P^{(n)}$. In particular, the set of Rees valuations of $P^{(n)}$ over all $n\geq 1$ is infinite.
The symbolic algebra $\oplus_{n\geq 0}P^{(n)}$ is not finitely generated, for instance by Corollary 5.2 [14], since the analytic spread of $P^{(n)}$ is $2=\dim R$ for all $n$.
We will show that the analytic spread
(48) $\ell(\oplus_{n\geq 0}P^{(n)})=\dim(\oplus_{n\geq 0}P^{(n)})/m_{R}(\oplus_{n\geq 0}P^{(n)})=0,$
where $m_{R}$ is the maximal ideal of $R$. Thus the fiber $\psi^{-1}(m_{R})=\emptyset$, where
$\psi:\mbox{Proj}(\oplus_{n\geq 0}P^{(n)})\rightarrow\mbox{Spec}(R)$
is the natural projection.
We now make the construction. Let $E$ be an elliptic curve over an algebraically closed field $k$ and $p\in E$ be a (closed) point. Let $\mathcal{L}=\mathcal{O}_{E}(p)$ and $V(\mathcal{L})=\mbox{Spec}(\oplus_{n\geq 0}\mathcal{L}^{n})$, the total space of the line bundle $\mathcal{L}^{-1}$, with canonical morphism $\pi:V(\mathcal{L})\rightarrow E$. By Propositions II.8.4.2 and II.4.4 [20], we have a natural inclusion
$V(\mathcal{L})\subset{\mathbb{P}}(\mathcal{L}\oplus\mathcal{O}_{E})\cong{\mathbb{P}}(\mathcal{O}_{E}\oplus\mathcal{L}^{-1})$
and the zero section $E_{0}$ of $\pi$ is the image of the inclusion ${\mathbb{P}}(\mathcal{L}^{-1})\subset{\mathbb{P}}(\mathcal{O}_{E}\oplus\mathcal{L}^{-1})$. By Proposition V.2.9 [22],
$\mathcal{O}_{V(\mathcal{L})}(E_{0})\otimes\mathcal{O}_{E_{0}}\cong\mathcal{O}_{{\mathbb{P}}(\mathcal{O}_{E}\oplus\mathcal{L}^{-1})}(E_{0})\otimes\mathcal{O}_{E_{0}}\cong\mathcal{O}_{E}(-p).$
Let $A=\oplus_{n\geq 0}\Gamma(E,\mathcal{L}^{n})$. $A$ is a normal domain. By Proposition II.8.8.2 and Remark II.8.8.3 [20], there exists a canonical projective morphism $V(\mathcal{L})\rightarrow\mbox{Spec}(A)$ which contracts the zero section of $E_{0}$ of $V(\mathcal{L})$ to the graded maximal ideal $\mathfrak{m}$ of $A$ and is an isomorphism on $V(\mathcal{L})\setminus E_{0}\rightarrow\mbox{Spec}(A)\setminus\{\mathfrak{m}\}$. Let $R=\hat{A}$, the completion of $A$ at $\mathfrak{m}$. $R$ is a complete local domain. Let $m_{R}$ be the maximal ideal of $R$. Let $X=V(\mathcal{L})\times_{\mbox{Spec}(A)}\mbox{Spec}(R)$, with natural birational projective morphism $\varphi:X\rightarrow\mbox{Spec}(R)$. We will identify $E_{0}$ with $E$.
There exists a point $q\in E$ such that $\mathcal{O}_{E}(q-p)$ has infinite order in the Jacobian of $E$. Since $R$ is complete, there exists a curve $C$ on the nonsingular scheme $X$ such that $C\cdot E=q$. Let $P=\Gamma(X,\mathcal{O}_{X}(-C))$, a height one prime ideal in $R$. The $n$-th symbolic power of $P$ is $P^{(n)}=\Gamma(X,\mathcal{O}_{X}(-nC))$. The Zariski decomposition of $C$ is $\Delta=C+E$ (Lemma 4.1 [11]) and $\Gamma(X,\mathcal{O}_{X}(-nC))=\Gamma(X,\mathcal{O}_{X}(-n\Delta))$ for all $n>0$ by Lemma 4.3 [11]. By Lemma 2.2 [11], for all $n>0$, the base locus of $-n\Delta$ is a subset of $E$. The curve $E$ is in the base locus of $-n\Delta$ for all $n>0$ since $\mathcal{O}_{X}(-n\Delta)\otimes\mathcal{O}_{E}\cong\mathcal{O}_{E}(n(q-p))$ and so $\Gamma(E,\mathcal{O}_{X}(-n\Delta)\otimes\mathcal{O}_{E})=0$. Thus
(49) $\Gamma(X,\mathcal{O}_{X}(-nC))=\Gamma(X,\mathcal{O}_{X}(-n\Delta-E))$
for all $n\geq 1$.
We have short exact sequences
(50) $0\rightarrow\mathcal{O}_{X}(-n\Delta-2E)\rightarrow\mathcal{O}_{X}(-n\Delta-E)\rightarrow\mathcal{O}_{X}(-n\Delta-E)\otimes\mathcal{O}_{E}\rightarrow 0$
By Theorem III.11.1 [22],
$H^{1}(X,\mathcal{O}_{X}(-n\Delta-2E))\cong\lim_{\leftarrow}H^{1}(X,\mathcal{O}_{X}(-n\Delta-2E)\otimes\mathcal{O}_{mE}),$
where the limit is over the schemes $mE$. By Serre duality on an elliptic curve, $H^{1}(E,\mathcal{O}_{E}(D))=0$ if $D$ is a divisor on $E$ of positive degree. From induction on $m$ in the short exact sequences
$0\rightarrow\mathcal{O}_{X}(-n\Delta-2E-mE)\otimes\mathcal{O}_{E}\rightarrow\mathcal{O}_{X}(-n\Delta-2E)\otimes\mathcal{O}_{(m+1)E}\rightarrow\mathcal{O}_{X}(-n\Delta-2E)\otimes\mathcal{O}_{mE}\rightarrow 0,$
we have that
$H^{1}(X,\mathcal{O}_{X}(-n\Delta-2E)\otimes\mathcal{O}_{(m+1)E})\cong H^{1}(X,\mathcal{O}_{X}(-n\Delta-2E)\otimes\mathcal{O}_{mE})$
for all $m\geq 1$. Since $H^{1}(X,\mathcal{O}_{X}(-n\Delta-2E)\otimes\mathcal{O}_{E})=0$, we have that
$H^{1}(X,\mathcal{O}_{X}(-n\Delta-2E)\otimes\mathcal{O}_{mE})=0$
for all $m\geq 1$. Thus
(51) $H^{1}(X,\mathcal{O}_{X}(-n\Delta-2E))=0$
for all $n$. By the Riemann Roch theorem on $E$, $h^{0}(\mathcal{O}_{E}(n(q-p)+p))=1$ for all $n$. Thus there exists a unique point $q_{n}\in E$ such that $\mathcal{O}_{E}(n(q-p)+p)\cong\mathcal{O}_{E}(q_{n})$. Since $q-p$ has infinite order in the Jacobian of $E$, we have that the $q_{n}$ are distinct points and $q_{n}\neq q$ for all $n$. By (50) and (51), we have surjections
$\Gamma(X,\mathcal{O}_{X}(-n\Delta-E))\rightarrow\Gamma(E,\mathcal{O}_{X}(-n\Delta-E)\otimes\mathcal{O}_{E})=\Gamma(E,\mathcal{O}_{E}(q_{n})).$
Thus the base locus of $\Gamma(X,\mathcal{O}_{X}(-n\Delta-E))$ is the point $q_{n}$, and so for all $n$,
(52) $P^{(n)}\mathcal{O}_{X}=\mathcal{O}_{X}(-n\Delta-E)\mathcal{I}_{n}$
where $\mathcal{I}_{n}$ is an ideal sheaf on $X$ such that the support of $\mathcal{O}_{X}/\mathcal{I}_{n}$ is $q_{n}$.
Let $\gamma_{n}:B_{n}\rightarrow\mbox{Spec}(R)$ be the normalization of the blow up of $P^{(n)}$. The Rees valuations of $P^{(n)}$ whose center on $R$ is $m_{R}$ are the valuations associated to the local rings $\mathcal{O}_{B_{n},G}$ where $G$ is a prime exceptional divisor of $\gamma_{n}$. Let $C_{n}$ be the normalization of the blowup of $P^{(n)}\mathcal{O}_{X}$. We have birational $\mbox{Spec}(R)$ morphisms
$\begin{array}[]{ccccc}&&C_{n}&&\\ &\swarrow&&\searrow&\\ B_{n}&&&&X.\\ \end{array}$
The morphism $C_{n}\rightarrow X$ is not an isomorphism and there does not exist an $R$-morphism $X\rightarrow B_{n}$ since $P^{(n)}\mathcal{O}_{X}$ is not invertible. By Zariski’s main theorem, there exists a curve $F$ on $C_{n}$ which maps to the point $q_{n}$ on $X$ and maps to a curve on $B_{n}$. Thus the valuation associated to the local ring $\mathcal{O}_{C_{n},F}$ is a Rees valuation of $P^{(n)}$. Since the points $q_{n}$ are all distinct, we conclude that the ideals $P^{(n)}$ have infinitely many Rees valuations.
It remains to show that if $m\neq n$, then there are no common Rees valuations of $P^{(m)}$ and $P^{(n)}$ which dominate the maximal ideal $m_{R}$ of $R$. By the above analysis, any exceptional curve for $B_{n}\rightarrow\mbox{Spec}(R)$ either comes from a curve on $C_{n}$ which contracts to $q_{n}$ on $X$ or is the image of the strict transform of $E$ on $C_{n}$. Since the $q_{n}$ are all distinct, we are reduced to showing that the valuation of $\mathcal{O}_{X,E}$ is not a Rees valuation of $P^{(n)}$ for any $n$.
Let $\alpha_{n}:Y_{n}\rightarrow X$ be the blow up of $q_{n}$. Let $C_{n}$ be the strict transform of $C$ on $Y_{n}$, $E_{n}$ be the strict transform of $E$ and let $F_{n}$ be the prime exceptional divisor of $Y_{n}\rightarrow X$. The curve $E_{n}$ is isomorphic to $E$. We have the intersection products $E_{n}\cdot C_{n}=q$ and $F_{n}\cdot E_{n}=q_{n}$. Further,
$-p=E\cdot E=E_{n}\cdot\alpha_{n}^{*}E=E_{n}\cdot(E_{n}+F_{n})=E_{n}\cdot E_{n}+q_{n}.$
Thus $E_{n}\cdot E_{n}=-p-q_{n}$, and so
(53) $(\alpha_{n}^{*}(-n\Delta-E)-F_{n})\cdot E_{n}=q_{n}-q_{n}=0.$
Since $\mathcal{O}_{X}/\mathcal{I}_{n}$ is supported at $q_{n}$, $\mathcal{I}_{n}\mathcal{O}_{Y_{n}}=\mathcal{O}_{Y_{n}}(-F_{n})\mathcal{J}_{n}$ where $\mathcal{J}_{n}$ is an ideal sheaf on $Y_{n}$ such that $\mathcal{O}_{Y_{n}}\mathcal{J}_{n}$ is supported on $F_{n}$. By (52), we have that
$P^{(n)}\mathcal{O}_{Y_{n}}=\mathcal{O}_{Y_{n}}(\alpha_{n}^{*}(-n\Delta-E)-F_{n})\mathcal{J}_{n}.$
Now
$\begin{array}[]{lll}P^{(n)}&=&\Gamma(Y_{n},\mathcal{O}_{Y_{n}}(\alpha_{n}^{*}(-n\Delta-E)-F_{n})\mathcal{J}_{n})\subset\Gamma(Y_{n},\mathcal{O}_{Y_{n}}(\alpha_{n}^{*}(-n\Delta-E)-F_{n}))\\ &\subset&\Gamma(Y_{n},\mathcal{O}_{Y_{n}}(-n\overline{C}))=P^{(n)},\end{array}$
so
$P^{(n)}=\Gamma(Y_{n},\mathcal{O}_{Y_{n}}(\alpha_{n}^{*}(-n\Delta-E)-F_{n})).$
Now $E_{n}$ is not in the base locus of $\alpha_{n}^{*}(-n\Delta-E)-F_{n}$ by (52), so the image of $\Gamma(Y_{n},\mathcal{O}_{Y_{n}}(\alpha_{n}^{*}(-n\Delta-E)-F_{n}))$ in $\Gamma(E_{n},\mathcal{O}_{Y_{n}}(-n\Delta-E)-F_{n})\otimes\mathcal{O}_{E_{n}})$ is nonzero. Further, $\mathcal{O}_{Y_{n}}(\alpha_{n}^{*}(-n\Delta-E)-F_{n})\otimes\mathcal{O}_{E_{n}}\cong\mathcal{O}_{E}$ by (53). Thus the image of $\Gamma(Y_{n},\mathcal{O}_{Y_{n}}(\alpha_{n}^{*}(-n\Delta-E)-F_{n}))$ is equal to $\Gamma(E_{n},\mathcal{O}_{E_{n}})=k$. Thus $E_{n}$ is disjoint from the base locus of $\alpha_{n}^{*}(-n\Delta-E)-F_{n}$, and thus $\mathcal{O}_{Y_{n}}/\mathcal{J}_{n}$ is supported at a finite number of points which are disjoint from $E_{n}$.
By continuing to blow up points above the support of $\mathcal{O}_{Y_{n}}/\mathcal{J}_{n}$, we construct a morphism $\beta_{n}:X_{n}\rightarrow Y_{n}$ such that $P^{(n)}\mathcal{O}_{X_{n}}$ is invertible. We then have that
$P^{(n)}\mathcal{O}_{X_{n}}=\mathcal{O}_{X_{n}}(-n\overline{C}_{n}-(n+1)\overline{E}_{n}-(n+2)\overline{F}_{n}-\sum a_{i}G_{i})$
where $\overline{C}_{n},\overline{E}_{n},\overline{F}_{n}$ are the respective strict transforms of $C_{n},E_{n}$ and $F_{n}$, $a_{i}$ are positive integers, and the $G_{i}$ are exceptional for $\beta_{n}$ and thus are disjoint from $\overline{E}_{n}$. We have that
$E_{n}\cdot(-n\overline{C}_{n}-(n+1)\overline{E}_{n}-(n+2)\overline{F}_{n}-\sum a_{i}G_{i})=0$
so the valuation associated to $\mathcal{O}_{X,E}=\mathcal{O}_{X_{n},\overline{E}_{n}}$ is not a Rees valuation of $P^{(n)}$ by Corollary 6.4 [11].
Finally, we prove equation (48). Let $Q=\oplus_{n\geq 0}\Gamma(X,\mathcal{O}_{X}(-n\Delta-E))$. $Q$ is a prime ideal in $\oplus_{n\geq 0}P^{(n)}=\oplus_{n\geq 0}\Gamma(X,\mathcal{O}_{X}(-n\Delta))$ (as explained before Remark 6.6 [11]) and since $E$ is the only exceptional curve of $\varphi:X\rightarrow\mbox{Spec}(R)$, $Q=\sqrt{m_{R}(\oplus_{n\geq 0}P^{(n)})}$ by Proposition 6.7 [11].
By Equation (49) and since $\Gamma(X,\mathcal{O}_{X}(-E))=\{f\in R\mid v_{E}(f)>0\}=m_{R}$, where $v_{E}$ is the valuation associated to $E$, $Q=m_{R}\oplus(\oplus_{n>0}P^{(n)})$. Thus
$\ell(\oplus_{n\geq 0}P^{(n)})=\dim\left[(\oplus_{n\geq 0}P^{(n)})/m_{R}(\oplus_{n\geq 0}P^{(n)})\right]=\dim\left[(\oplus_{n\geq 0}P^{(n)})/Q\right]=\dim R/m_{R}=0.$